### v2

# thanks: Equal contributionthanks: Equal contribution

Abstract
Ecological interactions can dramatically alter evolutionary outcomes in complex communities. Yet, the framework of population genetics largely neglects interactions from a species-rich community. Here, we bridge this gap by using dynamical mean-field theory to integrate community ecology into classical population genetics models. We show that ecological interactions result in emergent frequency-dependent selection between parents and mutants, characterized by a single parameter measuring the strength of ecological feedbacks. This result generalizes classical population genetics models to highly diverse communities and enables predictions of mutation outcomes in these eco-evolutionary settings. We derive an analytic expression for fixation probability that extends Kimura’s formula and reveals that ecological interactions strongly suppress the fixation of moderately beneficial mutations. This suppression arises because frequency-dependent selection leads to prolonged coexistence between parent and mutant lineages, which acts as a barrier to fixation. The strength of these effects increases with effective population size and the number of open niches in the ecosystem. Our study establishes a framework for integrating ecological interactions into population genetics, showing that evolutionary outcomes can be predicted using simple models even in the presence of complex community feedbacks.
preprint: MIT-CTP/5928
Introduction
A major goal of biology is to develop a quantitative theory of evolution. The mathematical framework of population genetics was a foundational step toward this goal. A cornerstone of population genetics is Kimura’s diffusion model (also called Wright-Fisher diffusion) [1, 2, 3, 4, 5], which conceptualizes the evolutionary dynamics of a mutant’s frequency as a stochastic process, subject to both diffusion (random genetic drift) and deterministic selection. Using this theory, one can predict two fundamental aspects of a mutant’s fate: its fixation probability and the time it takes to either fix in the population or go extinct.
Since then, it has been shown that a large class of evolutionary dynamics map onto Kimura’s diffusion model. Examples range from Wright-Fisher and Moran processes [4, 6] to single-species serial dilution experiments like Lenski’s long-term evolution [7, 8]. These developments have led to the widespread view that the results of classical population genetics are “universal” and apply broadly to evolving populations [4, 9, 10]. Indeed, subsequent studies have used Kimura’s basic framework as a starting point while extending evolution to other biological contexts, e.g., in fluctuating environments [11], and with spatial structure [12, 13].
While successful, these models make a simplifying assumption: that the evolutionary dynamics of mutants are mainly determined by their interaction with their parents. In doing so, they incorporate ecological context implicitly rather than explicitly. This makes it hard to understand how ecological interactions will alter evolutionary trajectories [14, 15, 16], which is important since almost all evolving populations are embedded in complex ecological communities with which they interact [17, 18, 19]. Any comprehensive theory of evolution must therefore explain how ecological interactions with a complex community will affect mutation outcomes. Without such an eco-evolutionary perspective, our understanding of evolution in natural contexts will remain incomplete [20, 21, 22].
Numerous works have sought to bridge this gap. These include population genetics models with frequency-dependent selection that capture more realistic evolutionary scenarios [23, 24, 25, 26]. However, since these models assume communities with one to two species, it is unclear whether these apply to more complex ecological settings [27, 28]. Other evolutionary theories seek to integrate ecological context [29, 30] but many of these efforts remain qualitative in nature and give rise to very different predictions [31, 32, 33, 34, 35, 36]. More quantitative approaches often make particular assumptions about the nature of ecological interactions that make it hard to assess their general applicability. This includes assuming low community diversity [37], specific fitness functions [38], and specific choices of ecological models [39, 40, 15]. In short, we still currently lack general understanding of how population genetics may predict and explain the dynamics of mutants in highly-diverse communities across a wide variety of ecological models.
At first, incorporating all the complex details of community ecological interactions might seem intractable due to the large number of unknown parameters. However, recent approaches from statistical physics and random matrix theory have revealed that for a sufficiently complex ecological community, it is possible to make strong predictions about ecological properties of an ecosystem such as its diversity or stability [41, 42, 43, 44, 45, 46, 47, 48]. A key insight from these works has been the realization that at the community level, the effects of ecological interactions can often be summarized in terms of a few emergent parameters that relate to susceptibilities that measure the strength of ecological feedback [49]. Inspired by these efforts, we may use tools from statistical physics to reveal links between ecological interactions and population genetics models.
Here we derive an effective theory of population genetics in complex ecological communities. We use dynamical mean-field theory (DMFT) to coarse-grain the effect of a highly-diverse community on a mutant’s frequency dynamics. Our central result is that community-mediated feedbacks generically lead to emergent frequency-dependent selection across a variety of ecological models. A community’s combined effect can be captured by only a single new parameter that quantifies the strength of community-mediated feedbacks on mutant dynamics. This parameter can be computed from a community’s statistical properties such as its niche-packing. As a result, population genetics models with frequency-dependent selection are effective in describing mutant dynamics also in complex communities. We further find that community effects dramatically alter the fixation probabilities of mutants compared to Kimura’s predictions. Further, in diverse communities, mutants can often coexist with their parents for extremely long periods, which drastically alters their fixation or extinction times. Our results highlight that the fate of a mutant can be quantitatively predicted by population genetics models even in complex ecological settings.
Figure 1: Population genetics and parent-mutant dynamics in the absence and presence of complex communities. (a) Classic population genetics typically models dynamics only for the parent and the mutant, independent of community context. (b) In complex communities, ecological interactions among species collectively influence the parent-mutant dynamics. As shown using dynamical mean-field theory, the ecological community mediates an emergent frequency-depdendent selection between the parent and the mutant. (c) Parent-mutant dynamics can be expressed in terms of mutant frequency $f$. (d) In classical population genetics, the dynamics of $f$ are driven by constant selection and stochastic drift. (e) In complex communities, additional frequency-dependent selection arises, with an emergent parameter $\eta$ characterizing the strength of ecological feedbacks. (f-i) Examples of dynamics of species abundances in classical population genetics (left) versus in complex communities (other species’ dynamics in grey). After a long time, the mutant eventually reaches either extinction (f–g) or fixation (h–i).
Results
Ecological feedbacks manifest as emergent frequency-dependent selection. Population genetics aims to describe how mutations spread through a species population. To do so, it follows the dynamics of a mutant’s frequency $f(t)$ in the population given by
$f(t)={N_{m}(t)\over N_{p}(t)+N_{m}(t)},$ (1)
where $N_{p}(t)$ and $N_{m}(t)$ are the abundances of parent and mutant strains at time $t$, respectively (Fig. 1a and c). For convenience, throughout this work, we use the term “classical population genetics” synonymously with the single-locus diffusion limit of population genetics in haploid populations [4]. Such models assume that mutant dynamics are shaped by two major evolutionary forces: (i) selection, which is characterized by a selection coefficient $s$ that measures the fitness difference between the mutant and parent, and (ii) stochastic drift, which is parameterized in terms of an effective population size $N_{\mathrm{eff}}$ (Fig. 1d). Stochasticity effects play an important role in mutational dynamics because mutants typically emerge at low frequencies $f(0)\ll 1$, making their early dynamics inherently noisy, e.g., due to randomness in birth and death events [4, 50]. At long times, mutants either fix ($f=1$) or go extinct ($f=0$, Fig. 1f and h).
Classical population genetics makes a simplifying but crucial assumption: the selection coefficient $s$ remains constant throughout the evolutionary process (Fig. 1d) [1, 2, 4], though see exceptions [11]. This assumption implies that ecological contribution in $s$, if any, remains constant. Effectively, this treats parents and mutants as if they are isolated from interactions with the dynamics of an ecological community. For this reason, it often misses the fact that ecological communities can potentially alter evolutionary trajectories through dynamical environmental feedbacks (Fig. 1b) [17, 18, 19]. This assumption becomes especially hard to justify in complex communities which are highly-diverse and ubiquitous in nature: from rainforests to microbiomes.
In this work we develop a general framework to understand how ecology affects mutant dynamics. Our central finding is that when mutations arise in a complex ecological community, the selection coefficient $s$ can no longer be treated as a constant or simple time-dependent function [11]. Instead, selection becomes frequency-dependent. This frequency-dependence emerges as a generic consequence of community-mediated ecological feedbacks on the parent and mutant (Fig. 1b).
To illustrate these ideas, we begin by analyzing the evolutionary dynamics of a mutant in an ecosystem modeled using a generalized Lotka-Volterra model (GLV) with demographic noise. The GLV model describes a complex, highly-diverse community of $S\gg 1$ species with abundances $N_{i}$ ($i=1,\ldots,S$) whose dynamics are governed by stochastic differential equations of the form
$\frac{dN_{i}}{dt}=N_{i}\left(r_{i}-N_{i}-\sum_{j\neq i,j=1}^{S}A_{ij}N_{j}\right)+\sqrt{2DN_{i}}\xi_{i}(t),$ (2)
with the $\{\xi_{i}(t)\}$ independent normal random variables with $\langle\xi_{i}(t)\rangle=0$ and $\langle\xi_{i}(t)\xi_{j}(t^{\prime})\rangle=\delta_{ij}\delta(t-t^{\prime})$. In this expression, $r_{i}$ is the carrying capacity of species $i$; $A_{ij}$ captures ecological interactions between species and encodes how species $j$ affects the abundance of species $i$; and $\sqrt{2DN_{i}}\xi_{i}(t)$ represents demographic noise. The parameter $D$ sets the overall scale of demographic noise and serves as an analog of a temperature or diffusion coefficient. Following recent work, we model parameters $r_{i}$ and $A_{ij}$ as random variables [43, 46, 51] drawn from Gaussian distributions (Methods), although our results remain valid for other distributions with finite mean and variance [52, 53].
To study population genetics in the GLV, we initialize a community with $S$ randomly chosen species at steady-state (Methods). We then pick one of these species at random to be the parent $p$. We denote the abundance of the parent species in the community by $N_{p}(t)$. At time $t=0$, we introduce a mutant $m$ to this community at a small abundance $N_{m}(0)\ll N_{p}(0)$. This is equivalent to assuming that the mutant starts at low frequency $f_{0}\ll 1$. To model the ecological similarity of parent and mutant, we assume that the mutant’s interaction parameters, $A_{mj}$, with other species in the community are correlated with those of the parent, $A_{pj}$. We denote the strength of this correlation as $\rho$, which is independent of any difference in carrying capacity between parent and mutant.
Consistent with population genetics, we assume that mutants and parents are highly correlated ($\rho\approx 1$) in their interactions with the community. With a high $\rho$, the total parent and mutant population behaves effectively as a single species from the community perspective, which crucially simplifies the ecological dynamics (SI Sec. F). Moreover, mutants and parents interact much more strongly with each other than with any other species in the community. To model this, we set parent-mutant interactions $A_{mp},A_{pm}$ to be close to but weaker than the intraspecific competition ($\approx 1$, SI Sec. A), while interactions with all other species in the community are assumed to be much weaker ($\approx 1/S$, SI Sec. A). However, as we will show below, even though mutants interact weakly with any individual species in the community, these weak effects add up due to the large number of species ($\approx S$) present in the community. Thus as a collective, community interactions can no longer be neglected and can strongly influence mutant dynamics. Throughout this work, we focus on the strong-selection-weak-mutation regime [54]: no subsequent mutations occur before a mutant fixes or goes extinct. For this reason, we focus on a single mutation and ignore the scenarios of multiple mutants (known as clonal interference) or mutation rates. We also assume that the community is at an ecological steady state when the mutant appears.
Our goal is to derive the effective dynamics of the mutant frequency $f(t)$ in a complex community described by the GLV. To do so, we use dynamical mean-field theory (DMFT) [42, 45, 41, 55]. The key idea behind DMFT is that in large, diverse communities the net effect of the community can be coarse-grained into an effective feedback on parent-mutant dynamics which is encoded in a single emergent parameter $\eta$ that summarizes the strength of ecological feedbacks (Fig. 1b). Using this technique, we find that mutant frequency can be described using the equation
$\frac{df}{dt}=\left(s_{\mathrm{inv}}-\eta f\right)f(1-f)+\sqrt{\frac{f(1-f)}{N_{\mathrm{eff}}}}\xi(t),$ (3)
where $s_{\mathrm{inv}}$ measures the invasion fitness of the mutant in the community. Since $s_{\mathrm{inv}}$ contains both the intrinsic selection coefficient $s$ (e.g., differences in carrying capacities) and community effects (SI Sec. A), it plays a role analogous to the mutant’s selective coefficient in classical population genetics. $N_{\text{eff}}=\left<N_{p}(0)+N_{m}(0)\right>/2D$ measures the effective population size of the combined parent-mutant population, which is approximately constant on average, especially for mutants where $\rho\approx 1$ (SI Sec. F). $\xi(t)$ is a random normal variable with $\langle\xi(t)\rangle=0$ and $\langle\xi(t)\xi(t^{\prime})\rangle=\delta(t-t^{\prime})$.
This equation results from two distinct eco-evolutionary processes: deterministic selection – the term proportional to f(1f)f(1-f) – and stochastic drift – the term proportional to $\xi(t)$. Remarkably, just introducing a single new parameter $\eta$ suffices to capture the collective effect of the entire community on parent-mutant dynamics. Eq. (3) shows that selection is now frequency-dependent since the selection coefficient $s_{\mathrm{inv}}-\eta f$ changes with mutant frequency. When $\eta=0$, e.g., in the absence of a community, we recover classical population genetics with a constant selection coefficient. However, when $\eta\neq 0$, ecological feedbacks fundamentally alter evolutionary dynamics.
[…]
Figure 2: Fixation probabilities suppressed by strong ecological effects. We compute fixation probabilities $p_{\mathrm{fix}}$ at various invasion fitness $s_{\mathrm{inv}}$ by simulating mutations within generalized Lotka-Volterra models. Error bars denote standard errors from multiple instances of demographic noise and mutants. (a) When ecological effects are weak relative to stochastic drift ($\alpha\ll 1$), our theory matches Kimura’s formula with $s=s_{\mathrm{inv}}$ for fixation probabilities and predicts them accurately in simulations. (b) When ecological effects are strong ($\alpha\gg 1$), Kimura’s formula applies only for deleterious or highly beneficial mutants. For moderately beneficial mutants with $s_{\mathrm{inv}}\lesssim\eta/2$, fixation probabilities are strongly suppressed relative to Kimura’s prediction. Our theory with ecology can capture such suppression and matches simulations. We give analytic expressions for fixation probability in (c) Kimura’s theory and (d) our theory with ecology. Ecological suppression of mutants depends on several factors: (e) For neutral mutations ($s_{\mathrm{inv}}=0$), we quantify the suppression by the percentage decrease in fixation probability from Kimura’s prediction ($1-p_{\mathrm{fix}}/f_{0}$). Suppression increases with larger effective population size $N_{\text{eff}}$ and lower parent-mutant correlation $\rho$. (f) Suppression is also stronger when the community is less packed (has more “open niches”), i.e., the number of surviving species is low relative to the bound set by competitive exclusion.
Generalization of Kimura’s formula for fixation probabilities. We now explore the consequences of community-mediated frequency-dependent selection on the fate of a mutant. One of the celebrated results of classical population genetics is Kimura’s formula for fixation probability $p_{\mathrm{fix}}$ [1], which describes how likely a mutant is to replace its parent and fix in a population (Fig. 2c). Kimura’s formula is given by
$p_{\mathrm{fix}}=\frac{1-\exp(-2N_{\mathrm{eff}}sf_{0})}{1-\exp(-2N_{\mathrm{eff}}s)}\,.$ (4)
This formula shows that as the fitness difference between mutant and parent $s$ increases, mutants become more likely to fix. Specifically, neutral mutants with $s=0$ fix with a probability equal to their initial frequency $p_{\mathrm{fix}}=f_{0}$. When selection is weak, i.e., $N_{\text{eff}}|s|\ll 1$, mutants are “effectively neutral” and fix with the same probability $f_{0}$. When selection is strong, i.e., $N_{\text{eff}}|s|\gg 1$, the formula reduces to $p_{\mathrm{fix}}=1-e^{-2N_{\text{eff}}sf_{0}}$ for positive $s$ and $p_{\mathrm{fix}}=e^{-2N_{\text{eff}}|s|}$ for negative $s$. Deleterious mutants with $s<0$ thus have a negligible fixation probability, while for beneficial mutants with $s>0$ the fixation probability increases linearly $p_{\mathrm{fix}}\propto s$. Ultimately, once a mutation is strongly beneficial ($s\gg 1/N_{\text{eff}}f_{0}$) the mutant will almost surely fix since $p_{\mathrm{fix}}\approx 1$.
Starting with Eq. (3), we solved the corresponding backward equation to obtain an analytic formula for the fixation probability $p_{\mathrm{fix}}$ in complex communities (SI Sec. D, see also [62]):
$p_{\mathrm{fix}}=\frac{\mathrm{Erfi}\left(\alpha\tilde{s}_{\mathrm{inv}}\right)-\mathrm{Erfi}\left(\alpha(\tilde{s}_{\mathrm{inv}}-f_{0})\right)}{\mathrm{Erfi}\left(\alpha\tilde{s}_{\mathrm{inv}}\right)-\mathrm{Erfi}\left(\alpha(\tilde{s}_{\mathrm{inv}}-1)\right)}\,$, (5)
where
$\mathrm{Erfi}(x)=\frac{2}{\sqrt{\pi}}\int_{0}^{x}dy\,e^{y^{2}}\,$, (6)
is the imaginary error function,
$\alpha=\sqrt{N_{\mathrm{eff}}\eta}\,$, (7)
is the ratio of the strength of ecological feedbacks and stochastic drift and
$\tilde{s}_{\mathrm{inv}}=\frac{s_{\mathrm{inv}}}{\eta}\,$, (8)
is the mutant invasion fitness normalized by the strength of the ecological feedback. The quantity $\tilde{s}_{\mathrm{inv}}$ is proportional to the “dressed invasion fitness” which was recently introduced in Ref. [14] for predicting the outcomes of ecological invasions. Eq. (5) serves as the generalization of Kimura’s formula in the context of complex ecological communities.
To gain intuition for this formula, it is useful to look at various limits of this expression. When ecological effects are weak, i.e., $\alpha\ll 1$, our formula reduces back to Kimura’s original formula in Eq. (4) (SI Sec. D) with $s=s_{\mathrm{inv}}$. To numerically confirm this, we simulated the dynamics of complex communities governed by the GLV model into each of which we introduced a mutant of a randomly chosen parent strain. We made ecological effects weak simply by setting a large value of $D$, increasing the strength of stochastic drift. We repeated simulations for mutants with a given invasion fitness $s_{\mathrm{inv}}$ and measured the fraction of simulations in which mutants fixed (Methods). As shown in Fig. 2a, the simulated fixation probabilities match our generalized formula remarkably well across a range of $s_{\mathrm{inv}}$. Further, in this limit, our formula virtually overlaps Kimura’s formula. In the next section, we will look at the opposite case of strong ecological effects, i.e., $\alpha\gg 1$, and show that we get a qualitatively different picture of the fate of a mutant compared to Kimura’s formula.
Ecological interactions strongly suppress mutant fixation. In the presence of strong ecological effects, i.e., when $\alpha\gg 1$, our numerical simulations for the fixation probability show a stark deviation from Kimura’s predictions (Fig. 2b), with an almost switch-like behavior with increasing $s_{\mathrm{inv}}$. This is in contrast with Kimura’s formula which shows a gradual, linear increase in fixation probability over a range of $s$. The deviation from Kimura is most pronounced for moderately beneficial mutants where ecological effects strongly suppress the fixation probability (Fig. 2b, blue region). The origin of this suppression is easiest to understand for neutral mutants ($s_{\mathrm{inv}}=0$) where our formula reduces to (SI Sec. D)
$p_{\mathrm{fix}}(s_{\mathrm{inv}}=0)=\begin{cases}f_{0}&\alpha\ll 1\\ e^{-\alpha^{2}}&\alpha\gg 1\end{cases}\,.$ (9)
This formula agrees with Kimura’s results when ecological effects are weak ($\alpha\ll 1$) and shows that ecological effects exponentially suppress neutral mutations when ecological effects are large ($\alpha\gg 1$).
Figure 3: Parent-mutant coexistence in complex communities. We compute mean absorption and extinction times at various invasion fitness $s_{\mathrm{inv}}$ by simulating mutations within generalized Lotka-Volterra models. Error bars denote standard errors from multiple instances of demographic noise. (a, b) The time scales of parent-mutant dynamics at various invasion fitness $s_{\mathrm{inv}}$ reveal the possibility of coexistence. Under strong ecological effects, the mean times to absorption (either extinction or fixation) (a) and to extinction (b) are exponentially longer for moderately beneficial mutants with $0\lesssim s_{\mathrm{inv}}\lesssim\eta$, peaking at $s_{\mathrm{inv}}\simeq\eta/2$ (dashed lines) as predicted by our theory. The maximum mean times also grow exponentially with the effective population size $N_{\mathrm{eff}}$. (c-h) Representative trajectories and phase portraits (in terms of the ratio between selection and drift) show distinct behaviors across different $s_{\mathrm{inv}}$. For low $s_{\mathrm{inv}}$ (f) or high $s_{\mathrm{inv}}$ (h), there is a single crossover (dashed lines) from drift to selection dominated regimes. Selection drives (green arrows) the mutant to extinction (c) or fixation (e). However, there are additional crossovers (g) when $s_{\mathrm{inv}}$ is moderate with $s_{\mathrm{inv}}\lesssim\eta$. Selection instead drives the mutant into a coexistence region centered at frequency $f^{*}$ (dark red), resulting in exponentially long times for parent-mutant coexistence (d).
Using Eq. (5), we can also compute the full range of invasion fitness $s_{\mathrm{inv}}$ for which mutants are suppressed when ecological effects are strong (SI Sec. D). We find that suppression occurs when
$0\lesssim s_{\mathrm{inv}}\lesssim s_{c}:=\frac{\eta}{2}\,$, (10)
where $s_{c}=\eta/2$ is the critical invasion fitness at which fixation probability switches from roughly 0 to almost 1. As we show in SI Sec. D, in this case the dominant contribution to ecological suppression in the range $0\lesssim s_{\mathrm{inv}}\lesssim s_{c}$ comes from the last term in the denominator of Eq. (5).
The ecological suppression of mutants is an emergent community-mediated phenomenon whose strength depends on $\alpha=\sqrt{N_{\text{eff}}\eta}$. To understand how different ecological properties of the underlying community control the strength of suppression, we measured $\alpha$ across a variety of parameter sets in the GLV model and plotted the corresponding deviation of the simulated fixation probability from Kimura’s predictions for a neutral mutation ($s_{\mathrm{inv}}=0$) (see Fig. 2e). Suppression increased with the effective population size $N_{\mathrm{eff}}$ (Fig. 2e). This can be understood by noting that increasing $N_{\mathrm{eff}}$ weakens stochastic drift, and hence increases the importance of ecological feedbacks. Suppression also increased with decreasing parent-mutant correlation $\rho$ (Fig. 2e). The reason for this is that the ecological feedback parameter $\eta$, and hence $\alpha$, is proportional to $1-\rho$ (SI Sec. AC). Intuitively, as parent and mutant become less similar (i.e., $\rho$ decreases), mutants interact less strongly with their parents, allowing community feedbacks to play a stronger role in the dynamics.
Most counterintuitively, suppression was stronger in less “packed” communities (Fig. 2f), where we have defined packing as the ratio of the number of non-extinct species to the maximum number of species allowed by competitive exclusion (often referred to as the species packing bound). This observation suggests that ecological feedbacks become weaker as communities get closer to a fully packed regime where all niches are occupied. To understand this counterintuitive effect, we draw on an analogy between ecological communities and mechanical jamming [63]. Just like jammed systems that become harder and harder to deform as they approach the jamming limit, packed communities are more “rigid” and less “deformable” as potential niches become occupied. In support of this analogy, we have verified that it is really packing that determines the strength of ecological feedbacks and that the number of surviving species by itself is not predictive of ecological suppression (SI Sec. A, B).
Finally, we show in the SI that the results in Fig. 2 generalize beyond the GLV model to many variants of consumer resource models. This leads us to conclude that ecological interactions likely generically suppress fixation of neutral to moderately beneficial mutants, strongly deviating from the predictions of classical population genetics.
Figure 4: Suppression of fixation probabilities explained by parent-mutant coexistence. (a) In classic population genetics, mutants are very likely to reach fixation after crossing the drift threshold (dashed gray line) at low frequencies, a process known as establishment. (b, c) With ecological effects, the coexistence region (dark red, Fig. 3(g)) modifies the fate of mutants. When the coexistence region is closer to extinction (b), i.e., $f^{*}<1/2$, the region acts as an additional barrier to fixation, causing extinction even after crossing the drift threshold. This mechanism accounts for the suppression of fixation probabilities (Fig. 3b). In contrast, the coexistence region no longer prevents mutants from fixation when it becomes closer to fixation (c), i.e., $f^{*}>1/2$.
Parent-mutant coexistence alters fixation and extinction times. In addition to fixation probability, the second important quantity in population genetics is the time it takes for a mutant to reach fixation or extinction (which we refer to as absorption time). This quantity characterizes the timescale of adaptation in evolutionary dynamics. It is also informative to study the establishment time or equivalently the time to extinction. These two time scales are intimately related since extinction probability becomes negligible as soon as the mutant becomes established in the population, long before fixation. The mean extinction time can also be theoretically estimated (SI Sec. H). These times are relatively short in classic population genetics. Namely, they reach a maximum of order $N_{\mathrm{eff}}f_{0}|\log f_{0}|$ [64] for neutral mutations ($s=0$), and decay with the magnitude of the fitness difference $|s|$.
On the other hand, we find that when ecological effects are strong ($\alpha\gg 1$), the mean times to absorption (Fig. 3a) and extinction (Fig. 3b) become exponentially large for moderately beneficial mutants (i.e., $s_{\mathrm{inv}}\lesssim\eta$). Both of these times show a sharp peak at the critical fitness difference $s_{\mathrm{inv}}\simeq\eta/2$ (Eq. (10)). This increase in absorption time is so dramatic that it can even be directly observed in simulation trajectories of mutant and parent dynamics (see Fig. 3c–e). We further find that the maximum of both the absorption and extinction times grows exponentially with the effective population size $N_{\mathrm{eff}}$. This is in contrast with Kimura’s theory where the maximum time grows linearly with $N_{\mathrm{eff}}$. The presence of these exponentially long time scales raises the possibility that parents and mutants can transiently coexist for long periods in complex communities, highlighting a crucial difference between classical population genetics and our eco-evolutionary approach.
Such coexistence is a direct consequence of frequency-dependent selection (Eq. (3)). To see this, it is helpful to think about the infinite effective population size limit of Eq. (3). In this limit, we can ignore stochasticity and focus on the fixed points of the deterministic dynamics which are given by the solutions to the equation $0=\left(s_{\mathrm{inv}}-\eta f\right)f(1-f)$. When the fitness satisfies $s_{\mathrm{inv}}>\eta$ or $s_{\mathrm{inv}}<0$, this equation has two fixed points at $f^{*}=0$ and $f^{*}=1$. On the other hand, when $0\leq s_{\mathrm{inv}}\leq\eta$, this equation has an additional fixed point at $f^{*}=\tilde{s}_{\mathrm{inv}}=s_{\mathrm{inv}}/\eta$, implying that the parent and mutant can potentially coexist in the absence of stochastic drift.
In the presence of drift, for small and large $\tilde{s}_{\mathrm{inv}}$ — i.e., for deleterious and strongly beneficial mutants — selection almost always drives mutants to extinction (Fig. 3c) and fixation (Fig. 3e) respectively. On the other hand, for moderately beneficial mutants the presence of the additional fixed point in the deterministic limit gives rise to additional crossovers (Fig. 3g) which create a “coexistence region” between extinction and fixation, centered at frequency $f^{*}\simeq\tilde{s}_{\mathrm{inv}}$. Selection drives the mutant into this region instead of extinction or fixation, hence the parent and mutant coexist for exponentially long times (Fig. 3d) before escaping the region (SI Sec. H). Such a region is absent in classical population genetics.
Parent-mutant coexistence also qualitatively explains the suppression of fixation probabilities for moderately beneficial mutants with $0<s_{\mathrm{inv}}<s_{c}$ (Fig. 2b). Recall that in classical population genetics, beneficial mutants reach fixation through establishment, i.e., crossing the drift threshold from drift to selection dominated regimes (Fig. 4a, see also Fig. 3f and h). The process of establishment is central to Kimura’s theory, since after this point, mutants grow almost deterministically towards fixation. However, this qualitative picture is modified by the coexistence region under strong ecological effects, as selection now drives mutants into this region after establishment (Fig. 3g). When the coexistence region is closer to extinction (Fig. 4b), i.e., $f^{*}<1/2$ or $s_{\mathrm{inv}}<s_{c}$, mutants that escape this region due to stochastic drift are more likely to go extinct than fix. Since such extinction occurs even after crossing the drift threshold, it suppresses fixation probabilities relative to Kimura’s prediction. On the other hand, when the coexistence region is closer to fixation (Fig. 4c), i.e., $f^{*}>1/2$ or $s_{\mathrm{inv}}>s_{c}$, it is more likely for mutants inside the region to drift to fixation instead, hence in this case coexistence does not suppress fixation.
Discussion
In this study, we developed an analytic framework for integrating ecological interactions with a complex community into population genetics models. Our calculations show that interactions with a highly-diverse community can dramatically affect evolutionary outcomes. Despite their complexity, ecological feedbacks give rise to a remarkably simple form of frequency-dependent selection between parent and mutant strains. Surprisingly, the entire community’s frequency-dependence can be captured by a single parameter $\eta$ which can be easily calculated from the community for a wide variety of ecological models. Thus our results reveal a principled method to apply existing population genetics models with frequency-dependent selection even in the presence of complex communities.
We find that when ecological feedback is strong, the fixation probabilities for moderately beneficial mutants are suppressed compared to Kimura’s predictions (Fig. 2b). The suppression increases with effective population size (Fig. 2e), weaker parent-mutant correlations (Fig. 2e), and decreases as ecosystems become more packed (Fig. 2f). We further show that the deviations from Kimura’s formula arise from prolonged parent-mutant coexistence arising from ecological interactions (Fig. 4b and c). This coexistence results in exponentially long absorption (Fig. 3a) and extinction times (Fig. 3b) for moderately beneficial mutants, as well as coexistence regions in the phase portraits of parent-mutant dynamics (Fig. 3g).
Our theory is robust to the details of the exact ecological model used to mathematically represent the community dynamics. We show that our framework applies to various ecological models including generalized Lotka-Volterra models and multiple variants of the consumer-resource model (SI Sec. AC). Moreover, our analysis suggests that parent-mutant coexistence is a generic feature of population genetics in complex communities and is ubiquitous across all of the ecological models we have studied. For this reason, we expect the framework developed here to be widely applicable.
Our theory makes testable predictions for high-resolution strain tracking in natural microbiomes [65, 66]. With high enough temporal resolution, strain dynamics should exhibit intermittent stochastic fluctuations in a range of frequencies above the drift barrier (Fig. 4b–c). These dynamics should be most pronounced for moderately beneficial mutants with $s_{\mathrm{inv}}\lesssim\eta/2$. Importantly, estimating both $f^{*}$ and $s_{\mathrm{inv}}$ from observed trajectories could enable direct inference of the ecological feedback strength $\eta$ from strain dynamics alone, suggesting an interesting new empirical method to quantify the impact of ecology during evolutionary dynamics [67].
Our findings also have broader implications for inference using population genetics models. Current methods largely neglect ecology and for this reason may yield biased estimates of effective population sizes, demographic histories, and selection coefficients. Designing predictors that incorporate $\eta$ into coalescent models and demographic inference frameworks has the potential to improve the accuracy of statistical methods, especially in settings where ecology plays an important role in the evolutionary dynamics. Our results also have important implications for conservation genetics. They suggest that traditional metrics that ignore ecology may under- or over-estimate extinction risks.
In the future, it will be useful to extend our theory to a wider range of ecological and evolutionary settings. On the ecological side, while our theory has so far focused on communities near steady state, ecosystems can exhibit more complex dynamical behavior such as limit cycles, multistability [68, 69], and chaos [42, 45, 70, 71, 72]. Extending our theory to these settings will enable us to understand how the fate of a mutant depends on the underlying dynamical behavior of a community. On the evolutionary side, for many species, especially microbes, mutation rates are sufficiently large so that multiple mutants can emerge before previous ones fix or go extinct [73]. In particular, new mutants can easily emerge when previous mutants are still coexisting transiently with their parents. Interactions between these mutants, known as clonal interference [74, 7], can significantly modify evolutionary dynamics. It will be useful to understand how such clonal interference affects the eco-evolutionary dynamics of complex ecological communities. Further, over longer timescales eco-evolutionary feedbacks may result in communities with structured ecological interactions compared with the randomly assembled communities we have focused on in this paper [14, 75, 76, 15, 39, 48]. It will also be useful to understand how to generalize our results to communities evolved over the long-term.
Methods
The details of all the analytical and numerical methods used can be found in the Supplementary Information. Briefly, we started with defining the equations for the generalized Lotka-Volterra model and various consumer-resource models, designating two of the species to be a parent and mutant. We analyzed these equations using dynamical mean-field theory and other approximations to derive Eq. (3) and the value of $\eta$. We then solved the resulting stochastic differential equation to obtain fixation probabilities (Eq. (5)), selection-drift ratio, and mean extinction time (Fig. 3). We performed numerical simulations by sampling random ecosystems at steady states, then explicitly solved the equations in the ecological models using the Euler method with demographic noise [77, 47, 78]. We repeated the same simulation with multiple instances of demographic noise to obtain the statistical quantities such as fixation probabilities and mean absorption/extinction times.
Acknowledgements. We would like to thank Michael M. Desai for helpful discussions. This work was funded by NIH NIGMS R35GM119461 to PM and Chan-Zuckerburg Initiative Investigator grant to PM. AG acknowledges support from the Ashok and Gita Vaish Junior Researcher Award, the DST-SERB Ramanujan Fellowship, as well the DAE, Govt. of India, under project no. RTI4001.
Supplementary Information
Appendix A Lotka-Volterra model
We begin with the details of our eco-evolution models. Consider an ecological community of $S$ competing species, with the abundance of each species $N_{i}$ (where $1\leq i\leq S$) described by the Lotka-Volterra dynamics:
$\frac{dN_{i}}{dt}=N_{i}\left(r_{i}-N_{i}-\sum_{j\neq i}A_{ij}N_{j}\right)+\sqrt{2DN_{i}}\xi_{i}(t)\,$, (11)
where $r_{i}$ is the carrying capacity of species $i$, and $A_{ij}$ is the competition coefficient between species $i$ and $j$. We have also introduced a white demographic noise $\xi_{i}(t)$ for each species with diffusion coefficient $D$, satisfying
$\left<\xi_{i}(t)\right>=0\,,\quad\left<\xi_{i}(t)\xi_{j}(t^{\prime})\right>=\delta_{ij}\delta(t-t^{\prime})\,.$ (12)
To model the diversity of this community, we work in the limit $S\gg 1$ and assume that $r_{i}$ and $A_{ij}$ are randomly drawn from Gaussian distributions, such that
$r_{i}=\mu_{r}+\sigma_{r}\delta r_{i}\,,\quad A_{ij}=\frac{\mu}{S}+\frac{\sigma}{\sqrt{S}}a_{ij}\,$, (13)
where $\mu_{r},\sigma_{r},\mu,\sigma$ are the distribution parameters, and $\delta r_{i},a_{ij}$ are zero-mean Gaussian random variables with
$\left[\delta r_{i}\delta r_{j}\right]=\delta_{ij}\,,\quad\left[a_{ij}a_{kl}\right]=\delta_{ik}\delta_{jl}+\gamma\delta_{il}\delta_{jk}\,$, (14)
where the square brackets denote averages over random parameters instead of the demographic noise. The reciprocity $1\leq\gamma\leq 1$ is the correlation between the off-diagonal entries $A_{ij}$ and $A_{ji}$. The results below also apply to other distributions with similarly defined mean and variance.
Suppose at time $t=0$, a mutation of a parent strain $p$ occurs and a new mutant strain $m$ with small abundance $N_{m}(0)\ll N_{p}(0)$ invades the community. The mutant is highly similar to its parent; we let the parent-mutant correlation of ecological interactions $\rho$ be close to $1$. We sample the competition coefficients for the mutant as
$a_{mi}$ $=\rho a_{pi}+\sqrt{1-\rho^{2}}z_{i}\,$, (15)
$a_{im}$ $=\rho a_{ip}+\sqrt{1-\rho^{2}}\left(\gamma z_{i}+\sqrt{1-\gamma^{2}}z_{i}^{\prime}\right)\,$, (16)
where $z_{i},z_{i}^{\prime}$ are independent zero mean, unit variance Gaussian random variables, and indices $i,j$ now denotes the rest of the community without the parent or the mutant. These choices of coefficients ensure the right correlations $\mathrm{Corr}\left(A_{pi},A_{mi}\right)=\mathrm{Corr}\left(A_{ip},A_{im}\right)=\rho$ and $\mathrm{Corr}\left(A_{mi},A_{im}\right)=\gamma$. We also have $\mathrm{Corr}\left(A_{pi},A_{im}\right)=\mathrm{Corr}\left(A_{mi},A_{ip}\right)=\gamma\rho$. On the other hand, the competition between the parent and the mutant should be much stronger than that with other species, controlled by their niche overlap. Therefore, the coefficients $A_{pm}$ and $A_{mp}$ should be of order $\rho\simeq 1$. In the main text, we assume
$A_{pm}=A_{mp}=\rho\,$, (17)
but here we leave them to be arbitrary for generality. As defined in Eq. (13), the other interaction coefficients $A_{pi},A_{mi},A_{ip},A_{im}$ (where $i\neq p,m$) are of order $1/S$ and much smaller than $A_{pm},A_{mp}$. Finally, we also allow arbitrary values for the carrying capacity $r_{m}$, in order to obtain the functional dependence of the fixation probability. In reality, we should also sample $r_{m}$ using $\rho$ similarly as above.
Including the parent-mutant dynamics, the full set of Lotka-Volterra equations are
$\frac{dN_{i}}{dt}$ $=N_{i}\left(r_{i}-N_{i}-\sum_{j\neq i}A_{ij}N_{j}-A_{ip}N_{p}-A_{im}N_{m}\right)+\sqrt{2DN_{i}}\xi_{i}(t)\,$, (18)
$\frac{dN_{p}}{dt}$ $=N_{p}\left(r_{p}-N_{p}-A_{pm}N_{m}-\sum_{i}A_{pi}N_{i}\right)+\sqrt{2DN_{p}}\xi_{p}(t)\,$, (19)
$\frac{dN_{m}}{dt}$ $=N_{m}\left(r_{m}-N_{m}-A_{mp}N_{p}-\sum_{i}A_{mi}N_{i}\right)+\sqrt{2DN_{m}}\xi_{m}(t)\,.$ (20)
We focus on the “strong-selection-weak-mutation” limit: we assume an ecological steady state and let the mutant invade when the community is close to the steady state, then we let the community evolve till the parent or the mutant becomes extinct, even though the parent and mutant may transiently coexist for a long time. This approximation is one of the main assumptions of our analysis, which ignores the possibility of having multiple mutations in the community simultaneously due to finite mutation rates.
Instead of the abundances themselves, it is useful to focus on the dynamics of the mutant frequency
$f=\frac{N_{m}}{N_{p}+N_{m}}\,$, (21)
which evolves till reaching the point $f=0$ or $f=1$. We denote the initial frequency as $f_{0}\ll 1$. Using the Lotka-Volterra equation, we have
$\frac{df}{dt}$ $=\frac{1}{(N_{p}+N_{m})^{2}}\left(N_{p}\frac{dN_{m}}{dt}-N_{m}\frac{dN_{p}}{dt}\right)$ (22)
$=f(1-f)\left(r_{m}-r_{p}-(1-A_{pm})N_{m}+(1-A_{mp})N_{p}-\sum_{i}(A_{mi}-A_{pi})N_{i}\right)+\sqrt{\frac{2Df(1-f)}{N_{p}+N_{m}}}\xi(t)\,.$ (23)
Here we have combined the noise terms of the parent and the mutant into a single white noise by adding their variances. The ecological feedback to the parent-mutant dynamics is determined by how the parent and the mutant impact other species in the community. We see that in Eq. (18), the abundances $N_{p},N_{m}$ serve only as a perturbation of order $1/S$. We can then model their influence to $N_{i}$ as a linear response. We write
$N_{i}(t)=N_{i/pm}(t)-\sum_{j}\int_{0}^{t}dt^{\prime}\,\nu_{ij}(t,t^{\prime})\left(A_{jp}N_{p}(t^{\prime})+A_{jm}N_{m}(t^{\prime})\right)\,$, (24)
where $N_{i/pm}$ is the abundance of species $i$ if both the parent and the mutant are excluded from the community, and
$\nu_{ij}(t,t^{\prime})=\frac{\partial N_{i}(t)}{\partial r_{j}(t^{\prime})}\,$, (25)
is the susceptibility kernel. Using the linear response, we can now write the community contribution in Eq. (22) as
$\sum_{i}(A_{mi}-A_{pi})N_{i}(t)=\sum_{i}(A_{mi}-A_{pi})N_{i/pm}(t)-\sum_{ij}\int_{0}^{t}dt^{\prime}\,\nu_{ij}(t,t^{\prime})(A_{mi}-A_{pi})\left(A_{jp}N_{p}(t^{\prime})+A_{jm}N_{m}(t^{\prime})\right)\,.$ (26)
At leading order, the last term can be simplified with the self-averaging property of the community:
$\sum_{ij}\int_{0}^{t}dt^{\prime}\,\nu_{ij}(t,t^{\prime})(A_{mi}-A_{pi})\left(A_{jp}N_{p}(t^{\prime})+A_{jm}N_{m}(t^{\prime})\right)$ (27)
$\simeq\,$ $\sum_{ij}\int_{0}^{t}dt^{\prime}\,\nu_{ij}(t,t^{\prime})\left[(A_{mi}-A_{pi})\left(A_{jp}N_{p}(t^{\prime})+A_{jm}N_{m}(t^{\prime})\right)\right]$ (28)
$\simeq\,$ $\frac{\gamma\sigma^{2}(1-\rho)}{S}\sum_{i}\int_{0}^{t}dt^{\prime}\,\nu_{ii}(t,t^{\prime})\left(N_{m}(t^{\prime})-N_{p}(t^{\prime})\right)$ (29)
$=\,$ $\gamma\sigma^{2}(1-\rho)\int_{0}^{t}dt^{\prime}\,\nu(t,t^{\prime})\left(N_{m}(t^{\prime})-N_{p}(t^{\prime})\right)\,$, (30)
where we have introduced the susceptibility
$\nu(t,t^{\prime})=\frac{1}{S}\sum_{i}\nu_{ii}(t,t^{\prime})\,.$ (31)
Therefore, the mutant frequency satisfies
$\frac{df}{dt}=\,$ $f(1-f)\left(r_{m}-r_{p}-(1-A_{pm})N_{m}+(1-A_{mp})N_{p}-\sum_{i}(A_{mi}-A_{pi})N_{i/pm}\right$. (32)
$\left.+\gamma\sigma^{2}(1-\rho)\int_{0}^{t}dt^{\prime}\,\nu(t,t^{\prime})\left(N_{m}(t^{\prime})-N_{p}(t^{\prime})\right)\right)+\sqrt{\frac{2Df(1-f)}{N_{p}+N_{m}}}\xi(t)\,.$ (33)
To obtain an equation involving $f$ only, further approximations to the abundances must be made. First, we may follow the traditional convention in population genetics to restrict the total parent-mutant abundance $N_{p}+N_{m}$ to be a constant. Such restriction is further justified when the parent-mutant correlation $\rho$ is high. As discussed in Sec. F, the total parent-mutant abundance varies little in average when $\rho$ is high and the whole community is near the ecological steady state. There are only fluctuations from demographic noise. Therefore, we can replace
$N_{p}(t)+N_{m}(t)\simeq\left<N_{p}(0)+N_{m}(0)\right>:=N_{0}\,$, (34)
and the effective number of individuals is $N_{\mathrm{eff}}=N_{0}/2D$. We recall that the angle bracket denotes the average over demographic fluctuations in the initial ecological steady state.
Next, since we start with the ecological steady state before the mutation, the community abundances $N_{i/pm}$ without the invasions of the parent and the mutant also vary little. These variations are further averaged out when $N_{i/pm}$ enters the sum in Eq. (32). Similar to the above, we can approximate
$\sum_{i}(A_{mi}-A_{pi})N_{i/pm}(t)\simeq\sum_{i}(A_{mi}-A_{pi})\left<N_{i/pm}(0)\right$>\,. (35)
As a remark, this approximation for $N_{i/pm}$ is different from the usual DMFT approximation. In the usual DMFT, the mean field $\sum A_{0i}N_{i/0}$ is treated as a colored noise, since the cavity abundance $N_{0}$ is stochastic representing the trajectories of all the abundances in the community. Here, the cavity abundances are only for the parent and the mutant instead of the rest of the community, hence we should take the corresponding realizations of the mean field instead of treating the mean field as noise.
Finally, although we do not have the full quantitative details of $\nu(t,t^{\prime})$, it is typically reasonable to assume the response to be much faster than the timescale of growth rates [41], especially when the community is near the ecological steady state. Empirically, we observe that to predict mutation outcomes, it suffices to approximate the susceptibility simply by $\nu(t,t^{\prime})\simeq\nu\delta(t-t^{\prime})$, where $\nu$ is the susceptibility of the steady state appeared in the static cavity method. As we will see, the approximation becomes less accurate when there are more indirect interactions in the ecological dynamics (Sec. E) or the parent-mutant correlation is not high (Sec. F). Quantifying the regimes where this approximation is controllable is left for future work.
With all these approximations, we arrive at
[…]
$\,\left.+2N_{0}\gamma\sigma^{2}\nu(1-\rho)(f-1/2)\right)+\sqrt{\frac{f(1-f)}{N_{\mathrm{eff}}}}\xi(t)$ (37)
$=$ $\,f(1-f)s(f)+\sqrt{\frac{f(1-f)}{N_{\mathrm{eff}}}}\xi(t)\,$, (38)
where $s(f)$ represents frequency-dependent selection. We see that $s(f)$ is linear in $f$. To better interpret the above result, we note that $s(f)$ is in fact the difference between the growth rates between the mutant and the parent at frequency $f$. The difference $s(f=0)$ is also the invasion fitness of the mutant $s_{\mathrm{inv}}$ and plays the same role as the selection coefficient $s$ in classic population genetics. Therefore, it is useful to rewrite $s(f)$ as a linear function in $f$ involving $s_{\mathrm{inv}}$:
$s(f)=s_{\mathrm{inv}}-\eta f\,$, (39)
where
$s_{\mathrm{inv}}=r_{m}-r_{p}+(1-A_{mp})N_{0}-\sum_{i}(A_{mi}-A_{pi})\left<N_{i/pm}(0)\right>-N_{0}\gamma\sigma^{2}\nu(1-\rho)\,$, (40)
and
$\eta=2N_{0}\left(1-\frac{A_{pm}+A_{mp}}{2}-\gamma\sigma^{2}\nu(1-\rho)\right)\,$, (41)
parametrizes the strength of the overall ecological effects. Note that the ecological feedback proportional to $\nu$ also contributes to the invasion fitness since the parent was already in the community before $t=0$. In conclusion, we get
$\frac{df}{dt}=f(1-f)\left(s_{\mathrm{inv}}-\eta f\right)+\sqrt{\frac{f(1-f)}{N_{\mathrm{eff}}}}\xi(t)\,$, (42)
which is the same dynamics as in the main text.
To better understand the impact of the community, we can estimate the susceptibility using the steady state condition for surviving species:
$r_{i}-\sum_{j}(I+A)^{*}_{ij}N_{j}=0\,$, (43)
where the “stars” means taking the components of surviving species only. The variations thus satisfies
$\delta N_{i}=(I+A)^{*-1}_{ij}\delta r_{j}\Rightarrow\nu=\frac{1}{S}\mathrm{Tr}(I+A)^{*-1}\,.$ (44)
The trace is simply the resolvent of $A^{*}$, which can be calculated using random matrix theory. The result is [79]
$\nu=\frac{1-\sqrt{1-4\gamma\sigma^{2}\phi}}{2\gamma\sigma^{2}}\,$, (45)
where $\phi$ is the fraction of surviving species. Since we have started with a near steady state of the community for the mutation, the above expression for $\nu$ should also hold approximately after the mutation. Therefore, we also have
$\eta=N_{0}\left(2-A_{pm}-A_{mp}-(1-\rho)\left(1-\sqrt{1-4\gamma\sigma^{2}\phi}\right)\right)\,.$ (46)
For the main text, we set $A_{pm}=A_{mp}=\rho$ and the above nicely factorizes into
$\eta=N_{0}(1-\rho)\left(1+\sqrt{1-4\gamma\sigma^{2}\phi}\right)\,.$ (47)
Note that this expression is valid only when $\phi<1/4\gamma\sigma^{2}$, which is an upper bound on the number of surviving species similar to May’s stability bound for ecological communities [44]. This is the species packing bound referred in the main text.
Appendix B MacArthur consumer-resource model
Following the same procedure as in Appendix A, it is straightforward to extend the results to consumer-resource models. We will see that the mutant frequency follows the same dynamics as before, only with a different $\eta$.
We first study the MacArthur consumer-resource model with self-renewing resources. We consider an ecological community of $S$ species and $M$ resources. The species abundances $N_{i}$ (where $1\leq i\leq S$) and resource abundances $R_{\alpha}$ (where $1\leq\alpha\leq M$) follow
$\frac{dN_{i}}{dt}$ $=N_{i}\left(\sum_{\beta}C_{i\beta}R_{\beta}-m_{i}\right)+\sqrt{2DN_{i}}\xi_{i}(t)\,$, (48)
$\frac{dR_{\alpha}}{dt}$ $=R_{\alpha}\left(K_{\alpha}-R_{\alpha}-\sum_{j}E_{j\alpha}N_{j}\right)\,$, (49)
where $K_{\alpha}$ is the carrying capacity of resource $\alpha$, $m_{i}$ is the intrinsic mortality rate for species $i$, $C_{i\alpha}$ represents the consumption preferences of species $i$ for resource $\alpha$, and $E_{i\alpha}$ is the corresponding impact of species $i$ on resource $\alpha$. The demographic noise $\xi_{i}(t)$ is the same as in the previous section. For a diverse community, we work in the limit $S,M\gg 1$ but finite $S/M$. We further assume that the parameters are randomly drawn from Gaussian distributions, such that
$K_{\alpha}=\mu_{K}+\sigma_{K}\delta K_{\alpha}\,,\quad m_{i}=\mu_{m}+\sigma_{m}\delta m_{i}\,$, (50)
$C_{i\alpha}=\frac{\mu}{M}+\frac{\sigma}{\sqrt{M}}c_{i\alpha}\,,\quad E_{i\alpha}=\frac{\mu}{M}+\frac{\sigma}{\sqrt{M}}e_{i\alpha}\,$, (51)
where $\mu_{K},\sigma_{K},\mu_{m},\sigma_{m},\mu,\sigma$ are the distribution parameters, and $\delta K_{\alpha},\delta m_{i},c_{i\alpha},e_{i\alpha}$ are zero-mean Gaussian random variables with
$\left[\delta K_{\alpha}\delta K_{\beta}\right]=\delta_{\alpha\beta}\,,\quad\left[\delta m_{i}\delta m_{j}\right]=\delta_{ij}\,$, (52)
$\left[c_{i\alpha}c_{j\beta}\right]=\left[e_{i\alpha}e_{j\beta}\right]=\delta_{ij}\delta_{\alpha\beta}\,,\quad\left[c_{i\alpha}e_{j\beta}\right]=\kappa\delta_{ij}\delta_{\alpha\beta}\,.$ (53)
The reciprocity $0\leq\kappa\leq 1$ of species-resource interaction is the correlation between the matrices $C_{i\alpha}$ and $E_{i\alpha}$.
Suppose at time $t=0$, a mutation of a parent strain $p$ occurs and a new mutant strain $m$ with small abundance $N_{m}(0)\ll N_{p}(0)$ invades the community. Let the parent-mutant correlation be $\rho$. We sample the new interaction coefficients as
$c_{m\alpha}$ $=\rho c_{p\alpha}+\sqrt{1-\rho^{2}}z_{\alpha}\,$, (54)
$e_{m\alpha}$ $=\rho e_{p\alpha}+\sqrt{1-\rho^{2}}\left(\kappa z_{\alpha}+\sqrt{1-\kappa^{2}}z_{\alpha}^{\prime}\right)\,$, (55)
where $z_{\alpha},z_{\alpha}^{\prime}$ are independent zero mean, unit variance Gaussian random variables. These choices of coefficients ensure the right correlation $\mathrm{Corr}(C_{p\alpha},C_{m\alpha})=\mathrm{Corr}(E_{p\alpha},E_{m\alpha})=\rho$ and $\mathrm{Corr}(C_{m\alpha},E_{m\alpha})=\kappa$. We leave the mortality rate $m_{m}$ to be arbitrary. Note that we do not introduce any new resources.
After introducing the mutant, the consumer-resource dynamics becomes
$\frac{dN_{p}}{dt}$ $=N_{p}\left(\sum_{\beta}C_{p\beta}R_{\beta}-m_{p}\right)+\sqrt{2DN_{p}}\xi_{p}(t)\,$, (56)
$\frac{dN_{m}}{dt}$ $=N_{m}\left(\sum_{\beta}C_{m\beta}R_{\beta}-m_{m}\right)+\sqrt{2DN_{m}}\xi_{m}(t)\,$, (57)
$\frac{dR_{\alpha}}{dt}$ $=R_{\alpha}\left(K_{\alpha}-R_{\alpha}-\sum_{j}E_{j\alpha}N_{j}-E_{p\alpha}N_{p}-E_{m\alpha}N_{m}\right)\,.$ (58)
In terms of the mutant frequency $f=N_{m}/(N_{p}+N_{m})$, the above becomes
$\frac{df}{dt}$ $=\frac{1}{(N_{p}+N_{m})^{2}}\left(N_{p}\frac{dN_{m}}{dt}-N_{m}\frac{dN_{p}}{dt}\right)$ (59)
$=f(1-f)\left(\sum_{\beta}(C_{m\beta}-C_{p\beta})R_{\beta}-m_{m}+m_{p}\right)+\sqrt{\frac{2Df(1-f)}{N_{p}+N_{m}}}\xi(t)\,.$ (60)
The dynamics of the other species $N_{i}$ will not be relevant in the following calculation. Now we adopt the same set of approximations as before. We can model the changes of $R_{\alpha}$ as linear responses:
$R_{\alpha}=R_{\alpha/pm}-\sum_{\beta}\int_{0}^{t}dt^{\prime}\,\chi_{\alpha\beta}(t,t^{\prime})\left(E_{p\beta}N_{p}(t^{\prime})+E_{m\beta}N_{m}(t^{\prime})\right)\,$, (61)
where
$\chi_{\alpha\beta}(t,t^{\prime})=\frac{\partial R_{\alpha}(t)}{\partial K_{\beta}(t^{\prime})}\,$, (62)
is the susceptibility kernel. Now we can write the resource contribution in Eq. (59) as
$\sum_{\alpha}(C_{m\alpha}-C_{p\alpha})R_{\alpha}(t)=\sum_{\alpha}(C_{m\alpha}-C_{p\alpha})R_{\alpha/pm}(t)-\sum_{\alpha\beta}\int_{0}^{t}dt^{\prime}\,\chi_{\alpha\beta}(t,t^{\prime})(C_{m\alpha}-C_{p\alpha})\left(E_{p\beta}N_{p}(t^{\prime})+E_{m\beta}N_{m}(t^{\prime})\right)$ (63)
We can then apply self-averaging to the products between $C$ and $E$ and approximate the last term as
$\sum_{\alpha\beta}\int_{0}^{t}dt^{\prime}\,\chi_{\alpha\beta}(t,t^{\prime})(C_{m\alpha}-C_{p\alpha})\left(E_{p\beta}N_{p}(t^{\prime})+E_{m\beta}N_{m}(t^{\prime})\right)$ (64)
$\simeq\,$ $\sum_{\alpha\beta}\int_{0}^{t}dt^{\prime}\,\chi_{\alpha\beta}(t,t^{\prime})\left[(C_{m\alpha}-C_{p\alpha})\left(E_{p\beta}N_{p}(t^{\prime})+E_{m\beta}N_{m}(t^{\prime})\right)\right]$ (65)
$\simeq\,$ $\frac{\kappa\sigma^{2}(1-\rho)}{M}\sum_{\alpha}\int_{0}^{t}dt^{\prime}\,\chi_{\alpha\alpha}(t,t^{\prime})\left(N_{m}(t^{\prime})-N_{p}(t^{\prime})\right)$ (66)
$=\,$ $\kappa\sigma^{2}(1-\rho)\int_{0}^{t}dt^{\prime}\,\chi(t,t^{\prime})\left(N_{m}(t^{\prime})-N_{p}(t^{\prime})\right)\,$, (67)
where we have introduced the susceptibility
$\chi(t,t^{\prime})=\frac{1}{M}\sum_{\alpha}\chi_{\alpha\alpha}(t,t^{\prime})\,.$ (68)
Therefore, the mutant frequency satisfies
$\frac{df}{dt}=\,$ $f(1-f)\left(\sum_{\beta}(C_{m\beta}-C_{p\beta})R_{\beta/pm}-m_{m}+m_{p}\right$. (69)
$\left.-\kappa\sigma^{2}(1-\rho)\int_{0}^{t}dt^{\prime}\,\chi(t,t^{\prime})\left(N_{m}(t^{\prime})-N_{p}(t^{\prime})\right)\right)+\sqrt{\frac{2Df(1-f)}{N_{p}+N_{m}}}\xi(t)\,.$ (70)
As before, we then further approximate
$\sum_{\beta}(C_{m\beta}-C_{p\beta})R_{\alpha/pm}(t)\simeq\sum_{\beta}(C_{m\beta}-C_{p\beta})\left<R_{\alpha/pm}(0)\right>\,,\quad N_{p}(t)+N_{m}(t)\simeq N_{0}\,,\quad\chi(t,t^{\prime})\simeq\chi\delta(t-t^{\prime})\,.$ (71)
We then arrive at:
$\frac{df}{dt}=$ $\,f(1-f)\left(\sum_{\beta}(C_{m\beta}-C_{p\beta})R_{\beta/pm}(0)-m_{m}+m_{p}\right$. (72)
$\,\left.-2\kappa\sigma^{2}\chi(1-\rho)N_{0}(f-1/2)\right)+\sqrt{\frac{f(1-f)}{N_{\mathrm{eff}}}}\xi(t)$ (73)
$=$ $\,f(1-f)s(f)+\sqrt{\frac{f(1-f)}{N_{\mathrm{eff}}}}\xi(t)\,.$ (74)
As before, we can separate the invasion fitness $s_{\mathrm{inv}}=s(f=0)$ from $s(f)$ and rewrite the dynamics in the form of Eq. (42) as
$\frac{df}{dt}=f(1-f)\left(s_{\mathrm{inv}}-\eta f\right)+\sqrt{\frac{f(1-f)}{N_{\mathrm{eff}}}}\xi(t)\,$, (75)
where the invasion fitness is
$s_{\mathrm{inv}}=\sum_{\beta}\left(C_{m\beta}-C_{p\beta}\right)\left<R_{\beta/pm}(0)\right>-m_{m}+m_{p}+\frac{\eta}{2}\,$, (76)
and the strength of ecological effect is
$\eta=2\kappa\sigma^{2}\chi(1-\rho)N_{0}\,.$ (77)
Now, we can estimate the susceptibility by the steady state conditions for surviving species and non-depleted resources (denoted by asterisks):
$\sum_{\beta}C^{*}_{i\beta}R_{\beta}-m_{i}=K_{\alpha}-R_{\alpha}-\sum_{j}E^{*}_{j\alpha}N_{j}=0\,.$ (78)
The variations satisfy
$\left(\begin{matrix}0&C^{*}\\ E^{*T}&I^{*}\end{matrix}\right)\left(\begin{matrix}\delta N\\ \delta R\end{matrix}\right)=\left(\begin{matrix}0\\ \delta K\end{matrix}\right)\,.$ (79)
Using block matrix inversion, we have
$\delta R=\left(I^{*}-E^{*T}\left(C^{*}E^{*T}\right)^{-1}C^{*}\right)\delta K\,$, (80)
hence,
$\chi$ $=\frac{1}{M}\mathrm{Tr}\left(I^{*}-E^{*T}\left(C^{*}E^{*T}\right)^{-1}C^{*}\right)$ (81)
$=\frac{1}{M}\mathrm{Tr}\,I^{*}-\frac{1}{M}\mathrm{Tr}\left(\left(C^{*}E^{*T}\right)^{-1}C^{*}E^{*T}\right)$ (82)
$=\phi_{R}-\frac{S}{M}\phi_{N}\,$, (83)
where $\phi_{R}$ is the fraction of non-depleted resources, and $\phi_{N}$ is the fraction of surviving species. In other words, $\chi$ measures how far the community is from the competitive exclusion limit. This $\chi$ is again the same as the corresponding susceptibility appeared in the static cavity method. We finally have
$\eta=2\kappa\sigma^{2}\left(\phi_{R}-\frac{S}{M}\phi_{N}\right)(1-\rho)N_{0}\,.$ (84)
Appendix C Generalized consumer-resource models
We can further extend the above results to more general consumer-resource models with abiotic resources. The dynamics take the form [49]
$\frac{dN_{i}}{dt}$ $=N_{i}\left(g_{i}(\vec{R})-m_{i}\right)+\sqrt{2DN_{i}}\xi_{i}(t)\,$, (85)
$\frac{dR_{\alpha}}{dt}$ $=h_{\alpha}(\vec{K})-q_{\alpha}(\vec{R})-f_{\alpha}(\vec{R},\vec{N})\,$, (86)
where $g_{i}(\vec{R})$ describes how the growth rate of species $i$ depends on resource abundances, $h_{\alpha}(\vec{K})$ is the external supply rate of resource $\alpha$ to the community, $q_{\alpha}(\vec{R})$ represents the resource dynamics in the absence of species, and $f_{\alpha}(\vec{R},\vec{N})$ is the rate at which resource $\alpha$ is produced or consumed by the species in the community. An example of such a model is the consumer-resource model with externally supplied resources:
$\frac{dN_{i}}{dt}$ $=N_{i}\left(\sum_{\alpha}C_{i\alpha}R_{\alpha}-m_{i}\right)+\sqrt{2DN_{i}}\xi_{i}(t)\,$, (87)
$\frac{dR_{\alpha}}{dt}$ $=K_{\alpha}-R_{\alpha}-\sum_{i}C_{i\alpha}N_{i}R_{\alpha}\,.$ (88)
The consumer preference matrix $C_{i\alpha}$ matches the definition in below. The sampling schemes for $C_{i\alpha},K_{\alpha},m_{i}$ are the same as in Sec. B.
Due to the generality, here we will derive the corresponding equation to Eq. (42) in terms of the functions $g_{i},h_{\alpha},q_{\alpha},f_{\alpha}$ themselves, but not the random matrix statistics. It will be helpful to define the following matrices [49]:
$C_{i\alpha}$ $=\left.\frac{\partial g_{i}}{\partial R_{\alpha}}\right|_{\vec{R}^{*}}\,$, (89)
$E_{i\alpha}$ $=\left.\frac{\partial f_{\alpha}}{\partial N_{i}}\right|_{\vec{R}^{*},\vec{N}^{*}}\,$, (90)
$Q_{\alpha\beta}$ $=\left.\frac{\partial f_{\alpha}}{\partial R_{\beta}}\right|_{\vec{R}^{*},\vec{N}^{*}}+\left.\frac{\partial q_{\alpha}}{\partial R_{\beta}}\right|_{\vec{R}^{*}}\,$, (91)
where $\vec{N}^{*},\vec{R}^{*}$ are the abundances at the steady state. $C_{i\alpha}$ and $E_{i\alpha}$ play similar roles as in the MacArthur consumer-resource model; see Eqs. (48) and (49). $Q_{\alpha\beta}$ represents the interactions between resources.
Suppose at time $t=0$, a mutation of a parent strain $p$ occurs and a new mutant strain $m$ with small abundance $N_{m}(0)\ll N_{p}(0)$ invades the community. The consumer-resource dynamics becomes
$\frac{dN_{p}}{dt}$ $=N_{p}\left(g_{p}(\vec{R})-m_{p}\right)+\sqrt{2DN_{p}}\xi_{p}(t)\,$, (92)
$\frac{dN_{m}}{dt}$ $=N_{m}\left(g_{m}(\vec{R})-m_{m}\right)+\sqrt{2DN_{m}}\xi_{m}(t)\,$, (93)
$\frac{dR_{\alpha}}{dt}$ $=h_{\alpha}(\vec{K})-q_{\alpha}(\vec{R})-f_{\alpha}(\vec{R},N_{i},N_{p},N_{m})$ (94)
$\simeq h_{\alpha}(\vec{K})-q_{\alpha}(\vec{R})-f_{\alpha}(\vec{R},N_{i})-E_{p\alpha}N_{p}-E_{m\alpha}N_{m}\,.$ (95)
In terms of the mutant frequency $f=N_{m}/(N_{p}+N_{m})$, the above becomes
$\frac{df}{dt}$ $=\frac{1}{(N_{p}+N_{m})^{2}}\left(N_{p}\frac{dN_{m}}{dt}-N_{m}\frac{dN_{p}}{dt}\right)$ (96)
$=f(1-f)\left(g_{m}(\vec{R})-g_{p}(\vec{R})-m_{m}+m_{p}\right)+\sqrt{\frac{2Df(1-f)}{N_{p}+N_{m}}}\xi(t)\,.$ (97)
The dynamics of the other species $N_{i}$ will not be relevant in the following calculation. Now we adopt the same set of approximations as before. We can model the changes of $R_{\alpha}$ as linear responses, which can be approximated by the variations of the steady state conditions:
$\left(\begin{matrix}0&C^{*}\\ E^{*T}&Q\end{matrix}\right)\left(\begin{matrix}\delta N\\ \delta R\end{matrix}\right)=\left(\begin{matrix}0\\ \delta h\end{matrix}\right)\,$, (98)
where
$\delta h_{\alpha}=-E_{p\alpha}N_{p}-E_{m\alpha}N_{m}\,.$ (99)
Using block matrix inversion, we have
$R_{\alpha}=R_{\alpha/pm}-\sum_{\beta}\left(\left(I-P\right)Q^{-1}\right)_{\alpha\beta}\left(E_{p\beta}N_{p}+E_{m\beta}N_{m}\right)\,$, (100)
where
$P=Q^{-1}E^{*T}\left(C^{*}Q^{-1}E^{*T}\right)^{-1}C^{*}\,.$ (101)
is a projector in resource space satisfying $P^{2}=P$ [49]. We see that the matrix $(I-P)Q^{-1}$ generalizes the susceptibility $\chi_{\alpha\beta}$ in the MacArthur consumer-resource model. Now, the mutant frequency satisfies
$\frac{df}{dt}=\,$ $f(1-f)\left(g_{m}\left(\vec{R}_{/pm}\right)-g_{p}\left(\vec{R}_{/pm}\right)-m_{m}+m_{p}\right$. (102)
$\left.-\sum_{\alpha\beta}(C_{m\alpha}-C_{p\alpha})\left(\left(I-P\right)Q^{-1}\right)_{\alpha\beta}\left(E_{p\beta}N_{p}+E_{m\beta}N_{m}\right)\right)+\sqrt{\frac{2Df(1-f)}{N_{p}+N_{m}}}\xi(t)\,.$ (103)
As before, we then approximate
$g_{m}\left(\vec{R}_{/pm}(t)\right)-g_{p}\left(\vec{R}_{/pm}(t)\right)\simeq\left<g_{m}\left(\vec{R}_{/pm}(0)\right)-g_{p}\left(\vec{R}_{/pm}(0)\right)\right>\,,\quad N_{p}(t)+N_{m}(t)\simeq N_{0}\,.$ (104)
We then arrive at the same equation as Eq. (42):
$\frac{df}{dt}=\,$ $f(1-f)\left(\left<g_{m}\left(\vec{R}_{/pm}(0)\right)-g_{p}\left(\vec{R}_{/pm}(0)\right)\right>-m_{m}+m_{p}\right$. (105)
$\left.-N_{0}\sum_{\alpha\beta}(C_{m\alpha}-C_{p\alpha})\left(\left(I-P\right)Q^{-1}\right)_{\alpha\beta}\left(E_{p\beta}(1-f)+E_{m\beta}f\right)\right)+\sqrt{\frac{f(1-f)}{N_{\mathrm{eff}}}}\xi(t)\,.$ (106)
$=$ $\,f(1-f)\left(s_{\mathrm{inv}}-\eta f\right)+\sqrt{\frac{f(1-f)}{N_{\mathrm{eff}}}}\xi(t)\,$, (107)
where the invasion fitness is
$s_{\mathrm{inv}}=$ $\,\left<g_{m}\left(\vec{R}_{/pm}(0)\right)-g_{p}\left(\vec{R}_{/pm}(0)\right)\right>-m_{m}+m_{p}$ (108)
$\,-N_{0}\sum_{\alpha\beta}(C_{m\alpha}-C_{p\alpha})\left(\left(I-P\right)Q^{-1}\right)_{\alpha\beta}E_{p\beta}$ (109)
and the strength of ecological effect is
$\eta=N_{0}\sum_{\alpha\beta}(C_{m\alpha}-C_{p\alpha})\left(\left(I-P\right)Q^{-1}\right)_{\alpha\beta}\left(E_{m\beta}-E_{p\beta}\right)\,.$ (110)
Here we do not specify the statistics of the matrices $C,E,Q$. Nevertheless, we expect that upon self-averaging, the expression of $\eta$ involves only the trace of $(I-P)Q^{-1}$ instead of the full matrix, corresponding to the scalar susceptibility $\chi$ in the MacArthur consumer-resource model. We then conclude that generalized consumer-resource models lead to the same dynamics of mutant frequency as in the MacArthur consumer-resource model.
Appendix D Derivation and analysis of fixation probability
From the previous sections, we have seen that all the ecological models give rise to the same dynamics of mutant frequency $f$, satisfying the stochastic differential equation
$\frac{df}{dt}=f(1-f)\left(s_{\mathrm{inv}}-\eta f\right)+\sqrt{\frac{f(1-f)}{N_{\mathrm{eff}}}}\xi(t)\,$, (111)
where $s_{\mathrm{inv}}$ is the invasion fitness, $\eta$ is the strength of ecological effects depending on details of the ecological model, and $N_{\mathrm{eff}}$ is the initial effective population size of the parent and the mutant.
The fixation probability can be obtained by solving the stochastic differential equation. Let $p(f,t|x)$ be the probability that the frequency is between $f$ and $f+df$ at time $t$ given that $f(0)=x$. The corresponding Kolmogorov backward equation is
$\frac{\partial p}{\partial t}=x(1-x)(s_{\mathrm{inv}}-\eta x)\frac{\partial p}{\partial x}+\frac{x(1-x)}{2N_{\mathrm{eff}}}\frac{\partial^{2}p}{\partial x^{2}}\,.$ (112)
Since $f$ stops at either 0 or $1$ after sufficiently long time, $p$ must approach the steady state $p^{*}(f,x)=p(f,\infty|x)$ given by
$0=x(1-x)(s_{\mathrm{inv}}-\eta x)\frac{\partial p^{*}}{\partial x}+\frac{x(1-x)}{2N_{\mathrm{eff}}}\frac{\partial^{2}p^{*}}{\partial x^{2}}\,$, (113)
$\Rightarrow\frac{\partial p^{*}}{\partial x}\propto\exp\left[N_{\mathrm{eff}}\left(\eta x^{2}-2s_{\mathrm{inv}}x\right)\right]\,.$ (114)
Using the stopping condition $p(1,0)=0$ and $p(1,1)=1$, we arrive at the fixation probability
$p_{\mathrm{fix}}$ $:=p(1,f_{0})=\frac{\int_{0}^{f_{0}}dx\,\partial p^{*}/\partial x}{\int_{0}^{1}dx\,\partial p^{*}/\partial x}$ (115)
$=\frac{\mathrm{Erfi}\left(\alpha\tilde{s}_{\mathrm{inv}}\right)-\mathrm{Erfi}\left(\alpha(\tilde{s}_{\mathrm{inv}}-f_{0})\right)}{\mathrm{Erfi}\left(\alpha\tilde{s}_{\mathrm{inv}}\right)-\mathrm{Erfi}\left(\alpha(\tilde{s}_{\mathrm{inv}}-1)\right)}\,$, (116)
where
$\mathrm{Erfi}(x)=\frac{2}{\sqrt{\pi}}\int_{0}^{x}dy\,e^{y^{2}}\,$, (117)
is the imaginary error function. We have denoted $\alpha=\sqrt{N_{\mathrm{eff}}\eta}$ as the ratio of strength between ecological effects and demographic noise, and $\tilde{s}_{\mathrm{inv}}=s_{\mathrm{inv}}/\eta$ is proportional to the “dressed invasion fitness” [14].
We can better understand the formula for $p_{\mathrm{fix}}$ using the asymptotic forms
$\mathrm{Erfi}(x)=\begin{cases}2x/\sqrt{\pi}&|x|\ll 1\\ e^{x^{2}}/x\sqrt{\pi}&|x|\gg 1\,.\end{cases}$ (118)
For example, when $s_{\mathrm{inv}}=0$,
$p_{\mathrm{fix}}(s_{\mathrm{inv}}=0)=\frac{\mathrm{Erfi}\left(\alpha f_{0}\right)}{\mathrm{Erfi}\left(\alpha\right)}\,.$ (119)
For $\alpha\ll 1$ (and $f_{0}\ll 1$ as we have assumed), we can use the linear approximation for all the terms, thus
$p_{\mathrm{fix}}(s_{\mathrm{inv}}=0,\alpha\ll 1)\simeq\frac{\alpha f_{0}}{\alpha}=f_{0}\,.$ (120)
For $\alpha\gg 1$, the denominator is always exponentially large and dominating. Ignoring the subdominating factors, we have
$p_{\mathrm{fix}}(s_{\mathrm{inv}}=0,\alpha\gg 1)\sim e^{-\alpha^{2}}\,.$ (121)
We see that $p_{\mathrm{fix}}(s_{\mathrm{inv}}=0)$ matches Kimura’s prediction when $\alpha\ll 1$, but exponentially suppressed by $\alpha$ otherwise.
The fixation probability for neutral mutations suggests that there is a crossover between the regimes of weak and strong ecological effects, characterized by small and large $\eta$ respectively. First, for small $\eta$ with $\alpha\ll 1$, we have $|\alpha\tilde{s}_{\mathrm{inv}}|\gg 1$ for any finite $s_{\mathrm{inv}}$ since $\alpha\tilde{s}_{\mathrm{inv}}$ scales as $\eta^{-1/2}$. Therefore, we can use the exponential approximation for all the terms and $p_{\mathrm{fix}}$ becomes
$p_{\mathrm{fix}}$ $\simeq\frac{\exp\left(\alpha^{2}\tilde{s}_{\mathrm{inv}}^{2}\right)-\exp\left(\alpha^{2}(\tilde{s}_{\mathrm{inv}}-f_{0})^{2}\right)}{\exp\left(\alpha^{2}\tilde{s}_{\mathrm{inv}}^{2}\right)-\exp\left(\alpha^{2}(\tilde{s}_{\mathrm{inv}}-1)^{2}\right)}$ (122)
$\simeq\frac{1-\exp(-2N_{\mathrm{eff}}s_{\mathrm{inv}}f_{0})}{1-\exp(-2N_{\mathrm{eff}}s_{\mathrm{inv}})}\,$, (123)
which is exactly Kimura’s formula for fixation probability with $s=s_{\mathrm{inv}}$. Combining the result for neutral mutation, we see that Kimura’s formula works equally well in ecological context as long as $\alpha\ll 1$. On the other hand, for large $\eta$ with $\alpha\gg 1$, the behavior of $p_{\mathrm{fix}}$ is divided into different regimes according to $s_{\mathrm{inv}}$. Now for $|\alpha\tilde{s}_{\mathrm{inv}}|\ll 1$, we again have
$p_{\mathrm{fix}}\simeq\frac{\mathrm{Erfi}\left(\alpha f_{0}\right)}{\mathrm{Erfi}\left(\alpha\right)}\sim e^{-\alpha^{2}}\,.$ (124)
For $|\alpha\tilde{s}_{\mathrm{inv}}|\gg 1$, we have
$p_{\mathrm{fix}}$ $\simeq\frac{\exp\left(\alpha^{2}\tilde{s}_{\mathrm{inv}}^{2}\right)-\mathrm{sgn}(\tilde{s}_{\mathrm{inv}}-f_{0})\exp\left(\alpha^{2}(\tilde{s}_{\mathrm{inv}}-f_{0})^{2}\right)}{\exp\left(\alpha^{2}\tilde{s}_{\mathrm{inv}}^{2}\right)-\mathrm{sgn}(\tilde{s}_{\mathrm{inv}}-1)\exp\left(\alpha^{2}(\tilde{s}_{\mathrm{inv}}-1)^{2}\right)}$ (125)
$=\frac{1-\mathrm{sgn}(\tilde{s}_{\mathrm{inv}}-f_{0})\exp\left(-\alpha^{2}f_{0}(2\tilde{s}_{\mathrm{inv}}-f_{0})\right)}{1-\mathrm{sgn}(\tilde{s}_{\mathrm{inv}}-1)\exp\left(-\alpha^{2}(2\tilde{s}_{\mathrm{inv}}-1)\right)}\,.$ (126)
The behavior of the above expression is further divided into two regimes. When $\tilde{s}_{\mathrm{inv}}\ll 1/2$, the denominator is dominating and $p_{\mathrm{fix}}$ is still exponentially suppressed:
$p_{\mathrm{fix}}\sim\exp\left(-2N_{\mathrm{eff}}(s_{\mathrm{inv}}-\eta/2)\right)\,.$ (127)
When $\tilde{s}_{\mathrm{inv}}\gg 1/2$, the denominator becomes very close to $1$ and the numerator becomes dominating instead. We then have
$p_{\mathrm{fix}}\simeq\frac{1-\exp(-2N_{\mathrm{eff}}s_{\mathrm{inv}}f_{0})}{1-\mathrm{sgn}(\tilde{s}_{\mathrm{inv}}-1)\exp(-2N_{\mathrm{eff}}s_{\mathrm{inv}})}\,$, (128)
which matches Kimura’s formula approximately for $\tilde{s}_{\mathrm{inv}}<1$ and exactly for $\tilde{s}_{\mathrm{inv}}>1$.
In conclusion, while $p_{\mathrm{fix}}$ for negative $s_{\mathrm{inv}}$ is always exponentially suppressed regardless of ecology, the presence of a large $\eta$ now also exponentially suppresses $p_{\mathrm{fix}}$ with positive $s_{\mathrm{inv}}$ up to
$s_{\mathrm{inv}}\simeq s_{c}:=\frac{\eta}{2}\,.$ (129)
We can then interpret $s_{c}$ as a new threshold in $s$ for beneficial mutations that overcome the competitive ecological effects. In particular, the result for neutral mutations can now be understood as
$p_{\mathrm{fix}}(s_{\mathrm{inv}}=0)\sim\exp(-2N_{\mathrm{eff}}s_{c})\,$, (130)
which is the same as Kimura’s prediction but for $s=-s_{c}$.
Appendix E Fixation probabilities in consumer-resource models
Figure S1: Theoretical and simulated fixation probabilities in consumer-resource models. We compute fixation probabilities $p_{\mathrm{fix}}$ at various invasion fitness $s_{\mathrm{inv}}$ by simulating mutations within consumer-resource models with (a) self-renewing resources (MacArthur consumer-resource model) and (b) externally supplied resources. In both cases, our theoretical predictions are more accurate than Kimura’s predictions, but do not fully fit with the simulation results. Error bars denote standard errors from multiple instances of demographic noise and mutants.
Figure S2: Ecological suppression of fixation probabilities controlled by species packing in consumer-resource models. We compute fixation probabilities $p_{\mathrm{fix}}$ at various invasion fitness $s_{\mathrm{inv}}$ by simulating mutations within MacArthur consumer-resource models with different packing fractions $S^{*}/M^{*}$. The ecological suppression of fixation probabilities is stronger when the packing fraction is lower. Error bars denote standard errors from multiple instances of demographic noise and mutants.
In the main text, we have compared our predictions on $p_{\mathrm{fix}}$ with the simulation results from generalized Lotka-Volterra model. The comparison can similarly be done for various consumer-resource models as in Figure S1. We find that our predictions are in general more accurate than Kimura’s predictions. On the other hand, there are some deviations for $0<s_{\mathrm{inv}}<\eta$, where the parent and the mutant can transiently coexist, hence the deterministic ecological dynamics affect the fixation probabilities more significantly. Compared to generalized Lotka-Volterra model, here the species interactions are mediated through resource dynamics. For self-renewing resources, there are additional complications since resources can also go extinct. We suspect that these additional complexities cause our DMFT approximations in Sec. B and C to be less accurate.
We have also demonstrated in the main text that the ecological suppression of fixation probabilities is affected by species packing. While the packing is related to May’s stability bound for generalized Lotka-Volterra model, for consumer-resource models the packing is related to the competitive exclusion principle. The packing fraction is given by $S^{*}/M^{*}$, where $S^{*}$ is the number of surviving species and $M^{*}$ is the number of non-depleted resources. For externally supplied resources we have $M^{*}=M$. The packing bound is $S^{*}/M^{*}\leq 1$ for self-renewing resources and interestingly $S^{*}/M^{*}\leq 1/2$ for externally supplied resources [63]. In Fig. S2, we indeed see that the fixation probabilities are more suppressed for a less packed community with lower packing fraction.
In conclusion, we see that the results for different consumer-resource models are qualitatively the same as in generalized Lotka-Volterra model. Such agreement suggests that our theory is indeed applicable to various ecological models.
Appendix F Effects of parent-mutant correlation on the mutation dynamics
Figure S3: Dynamics of parent and mutant abundances during fixation of a neutral mutation at various parent-mutant correlation. Fixation of a neutral mutation requires deviation from the deterministic ecological steady state. At high $\rho=0.9$ (a), the total parent-mutant abundance remains constant on average throughout the full dynamics. At low $\rho=0.5$ (b), the total parent-mutant abundance drops significantly towards the final mutant abundance.
Figure S4: Theoretical and simulated fixation probabilities in generalized Lotka-Volterra models at various parent-mutant correlation. We compute fixation probabilities $p_{\mathrm{fix}}$ at various invasion fitness $s_{\mathrm{inv}}$ by simulating mutations within generalized Lotka-Volterra models. Our theory (solid curves) matches the simulations (dots) for high $\rho$, but deviates from the simulations for low $\rho$ (a). The relative error is more significant for lower $s_{\mathrm{inv}}$ (b). Error bars denote standard errors from multiple instances of demographic noise and mutants.
Throughout the paper, we have assumed high parent-mutant correlation $\rho\approx 1$. In fact, this is one of the crucial assumptions that allows us to simplify the mutation dynamics into the form in Eq. (42). Here we evaluate in more detail how parent-mutant correlation affects the mutation dynamics.
Qualitatively, when the parent and the mutant are highly similar in their ecological interactions with the community, the total parent-mutant population behaves effectively as a single species from the community perspective. Therefore, the community as a whole stays near the ecological steady state. This is true even when the parent and the mutant deviate much away from their steady states due to demographic noise; such deviation is required for a mutant with low $s_{\mathrm{inv}}$ to fix and replace the parent.
The fixation dynamics simplifies in the vicinity of the community steady state in the following ways. First, the total parent-mutant abundance $N_{p}+N_{m}$ always stays constant on average, with fluctuations coming from demographic noise only. In contrast, when $\rho$ is not high, $N_{p}+N_{m}$ is not stationary on average when a mutant with low $s$ fixes, as shown in Fig. S3. Second, the community response throughout the dynamics can be approximated by the one at the steady state. Correspondingly, the susceptibility $\nu(t,t^{\prime})$ can be replaced by $\nu\delta(t-t^{\prime})$ as predicted using the cavity method for the steady state. As shown in Fig. S4, this approximation for the DMFT indeed becomes less accurate when $\rho$ is not high, leading to errors in fixation probabilities. We suspect that a more careful treatment of the DMFT can predict mutation outcomes without high $\rho$.
Appendix G Comparison between selection and drift dynamics
In this appendix, we obtain a qualitative understanding of both the results in classic population genetics and our new results, by comparing the relative sizes of selection and drift terms in the differential equation for $f$. This analysis helps us understand how parent-mutant coexistence arises in the presence of ecological effects.
G.1 Without ecological effects
For pedagogical purposes, it is helpful to start by re-deriving some classic results from population genetics in the absence of ecology. As explained in the main text, a mutant starts to invade at a very small frequency where genetic drift dominates. The mutant can either become extinct or grow to a larger frequency due to drift. This process continues until the surviving mutants reach a frequency $f\sim 1/N_{\mathrm{eff}}|s|$ , where selection becomes dominant instead. This critical frequency is the drift threshold as shown in Fig. 4a in the main text.
We now obtain a heuristic analytical estimate of drift threshold, following the approach in [80, 81]. Recall that the dynamics of $f$ in classic population genetics is given by
$\frac{df}{dt}=f(1-f)s+\sqrt{\frac{f(1-f)}{N_{\mathrm{eff}}}}\xi(t)\,.$ (131)
The exact solution to this differential equation is in general complicated as both the growth rate and diffusion rate depends on $f$. On the other hand, analogous to the Euler method for numerically solving differential equations, we can approximate the dynamics of $f$ by splitting it into infinitesimal time intervals with width $\Delta t$. Within a time interval around some time $t$, the changes in both the growth rate and the diffusion rate can be neglected. Therefore, the change in frequency $\Delta f$ within the time interval can be easily solved and is given by
$\Delta f:=f(t+\Delta t)-f(t)\simeq f(t)(1-f(t))s\Delta t+\sqrt{\frac{f(t)(1-f(t))\Delta t}{N_{\mathrm{eff}}}}z\,$, (132)
where $z$ is a zero mean, unit variance Gaussian random variable. Below, we can simply use $f$ to denote $f(t)$.
Let us focus on small frequencies $f\ll 1$ where drift dominates over selection. In this case, $\Delta f$ can be either positive or negative due to randomness, and its magnitude can be estimated by the variance
$|\Delta f|\sim\sqrt{\frac{f(1-f)\Delta t}{N_{\mathrm{eff}}}}\,.$ (133)
The above approximation of constant growth and diffusion rates remains valid until $\Delta f$ is comparable to $f$, that is
$|\Delta f|\sim\sqrt{\frac{f(1-f)\Delta t}{N_{\mathrm{eff}}}}\sim f\Rightarrow\Delta t\sim\frac{N_{\mathrm{eff}}f}{1-f}\,.$ (134)
Within this time interval, the contribution to $\Delta f$ from selection is
$|\Delta f|_{\mathrm{sel}}\sim f(1-f)|s|\Delta t\sim|s|N_{\mathrm{eff}}f^{2}\,.$ (135)
Note that the absolute value is needed since $\Delta f$ due to selection can also be either positive or negative depending on the sign of $s$. To ensure that the approximation is self-consistent, $\Delta f$ due to selection must be negligible, i.e., $|\Delta f|_{\mathrm{sel}}\ll|\Delta f|$. Therefore, we require
$|s|N_{\mathrm{eff}}f^{2}\ll f\Rightarrow f\ll\frac{1}{N_{\mathrm{eff}}|s|}\,.$ (136)
Indeed, we see that drift dominates the dynamics when $f$ is below drift threshold.
G.2 With ecological effects
We can repeat the above analysis under the presence of ecological effects. As explained in the main text, when $0\lesssim s_{\mathrm{inv}}\lesssim\eta$, there is an additional deterministic fixed point $f^{*}\simeq\tilde{s}_{\mathrm{inv}}$. We then expect a new region around both sides of $f^{*}$ where drift dominates the dynamics, apart from the one near $f=0$. This is because at every fixed point, by definition, the deterministic contribution to the dynamics vanishes.
Suppose the mutant reaches frequency $f\gg f_{0}$ at time $t$. Again, we focus on an infinitesimal time interval with width $\Delta t$. From Eq. (42), we can approximate $\Delta f$ within this time interval as
$\Delta f\simeq f(1-f)(s_{\mathrm{inv}}-\eta f)\Delta t+\sqrt{\frac{f(1-f)\Delta t}{N_{\mathrm{eff}}}}z\,$, (137)
where $z$ is a zero mean, unit variance Gaussian random variable. Note that the variance is the same as in the case without ecological effects. Therefore, if drift dominates the dynamics, the above expression of $\Delta f$ remains a good approximation until
$\Delta t\sim\frac{N_{\mathrm{eff}}f}{1-f}\,$, (138)
which is the same as Eq. (134). To ensure that this approximation is self-consistent, the contribution to $\Delta f$ from selection must be negligible. Therefore, we require
$\left|f(1-f)(s_{\mathrm{inv}}-\eta f)\Delta t\right|\sim|s_{\mathrm{inv}}-\eta f|N_{\mathrm{eff}}f^{2}\ll f\,.$ (139)
Note again that the absolute value is needed since frequency can either increase or decrease, i.e., $\Delta f$ due to selection can be either positive or negative, particularly when $f$ is around the new fixed point $f^{*}\simeq\tilde{s}_{\mathrm{inv}}$, which is between $f=0$ and $f=1$. Based on the above condition, we can define the selection-drift ratio
$R=|s_{\mathrm{inv}}-\eta f|N_{\mathrm{eff}}f\,$, (140)
such that drift (selection) dominates the dynamics when $R<1$ ($R>1$), and $R=1$ is the crossover boundary between selection and drift dominated regimes. This is the ratio as shown in Fig. 3f-h in the main text.
Figure S5: Phase portraits of frequency dynamics in terms of selection-drift ratio. The selection-drift ratios $R$ (Eq. (140)) are calculated in different regimes of invasion fitness $s_{\mathrm{inv}}$, namely (a) $s_{\mathrm{inv}}<\sqrt{4\eta/N_{\mathrm{eff}}}$, (b) $\sqrt{4\eta/N_{\mathrm{eff}}}<s_{\mathrm{inv}}<\eta$, and (c) $s_{\mathrm{inv}}>\eta$. When $s_{\mathrm{inv}}$ is in the intermediate range, there are more crossovers between selection and drift dominated regimes, hence a coexistence region centered at $f^{*}$ emerges. The selection-drift ratios are identical to the ones in Fig. 3 in the main text.
Since we already know that our results reduce back to those in population genetics when $\alpha\ll 1$ ($\eta\approx 0)$, we now focus on the case of large $\eta$ and $\alpha\gg 1$. In this regime, due to the absolute value in $R$, the solutions to $R=1$ have different types of behavior depending on the value of $s$. Recall that we require $0\leq f\leq 1$. When $s_{\mathrm{inv}}<\sqrt{4\eta/N_{\mathrm{eff}}}$ (or $\tilde{s}_{\mathrm{inv}}<2/\alpha$), there is only one solution with $s_{\mathrm{inv}}-\eta f<0$: (Fig. S5a)
$f_{\mathrm{barrier}}\sim\frac{2}{N_{\mathrm{eff}}|s_{\mathrm{inv}}|}\left(\sqrt{1+\frac{4\eta}{N_{\mathrm{eff}}s_{\mathrm{inv}}^{2}}}-\mathrm{sgn}(s_{\mathrm{inv}})\right)^{-1}\,.$ (141)
When $s_{\mathrm{inv}}>\eta$, there is also only one solution but with $s_{\mathrm{inv}}-\eta f>0$: (Fig. S5c)
$f_{\mathrm{barrier}}\sim\frac{2}{N_{\mathrm{eff}}s_{\mathrm{inv}}}\left(\sqrt{1-\frac{4\eta}{N_{\mathrm{eff}}s_{\mathrm{inv}}^{2}}}+1\right)^{-1}\,.$ (142)
Therefore, the dynamics of $f$ is qualitatively the same as the one in classic population genetics for $s_{\mathrm{inv}}<\sqrt{4\eta/N_{\mathrm{eff}}}$ or $s_{\mathrm{inv}}>\eta$, with $f_{\mathrm{barrier}}$ being the drift threshold.
In contrast, the dynamics of $f$ is qualitatively distinct for $\sqrt{4\eta/N_{\mathrm{eff}}}<s_{\mathrm{inv}}\lesssim\eta$. We see that there are three solutions to $R=1$: (Fig. S5b)
$f_{\mathrm{barrier}}$ $=\frac{2}{N_{\mathrm{eff}}s_{\mathrm{inv}}}\left(1+\sqrt{1-\frac{4\eta}{N_{\mathrm{eff}}s_{\mathrm{inv}}^{2}}}\right)^{-1},$ (143)
$f_{\mathrm{coexist}}^{\mathrm{lower}}$ $=\frac{2}{N_{\mathrm{eff}}s_{\mathrm{inv}}}\left(1-\sqrt{1-\frac{4\eta}{N_{\mathrm{eff}}s_{\mathrm{inv}}^{2}}}\right)^{-1},$ (144)
$f_{\mathrm{coexist}}^{\mathrm{upper}}$ $=\frac{2}{N_{\mathrm{eff}}s_{\mathrm{inv}}}\left(\sqrt{1+\frac{4\eta}{N_{\mathrm{eff}}s_{\mathrm{inv}}^{2}}}-1\right)^{-1}\,$, (145)
with $f_{\mathrm{barrier}}<f_{\mathrm{coexist}}^{\mathrm{lower}}<f_{\mathrm{coexist}}^{\mathrm{upper}}$. Note that we have $f_{\mathrm{coexist}}^{\mathrm{lower}}\simeq f_{\mathrm{coexist}}^{\mathrm{upper}}\simeq f^{*}\simeq\tilde{s}_{\mathrm{inv}}$, hence these two frequencies correspond to the bounds of the coexistence region centered at $f^{*}$.
As demonstrated in the main text, our results in fixation probability and mean absorption/extinction time can be qualitatively understood using the nonlinear behavior of $R$ (Fig. S5b). Suppose a mutant starts with $f=f_{0}\ll 1$ at the beginning. It must overcome the first barrier at $f=f_{\mathrm{barrier}}$ or it becomes extinct. After that, selection brings the mutant to the region $f_{\mathrm{coexist}}^{\mathrm{lower}}<f<f_{\mathrm{coexist}}^{\mathrm{upper}}$, which is the coexistence region. Note that the deterministic fixed point $f^{*}\simeq\tilde{s}_{\mathrm{inv}}$ is within this region. Since $\left<df/dt\right>>0$ when $f<f_{\mathrm{coexist}}^{\mathrm{lower}}$ and $\left<df/dt\right><0$ when $f>f_{\mathrm{coexist}}^{\mathrm{upper}}$, both boundaries of the coexistence region are reflecting and the mutant is trapped inside this region. It can escape this region and reach near the absorbing boundaries $f=0,1$ only when the demographic noise reaches the tail of its distribution with exponentially low probability.
Now we see that if $s_{\mathrm{inv}}<s_{c}=\eta/2$, the coexistence region is closer to $f=0$ and it is exponentially more likely for the demographic noise to drive the mutant to extinction. Therefore, $p_{\mathrm{fix}}$ is exponentially suppressed for $s_{\mathrm{inv}}<s_{c}$. In contrast, when $s_{\mathrm{inv}}>s_{c}$, the coexistence region is closer to $f=1$ and it is exponentially more likely for the demographic noise to drive the mutant to fixation. The mutant becomes extinct mostly because it cannot overcome the first barrier at $f=f_{\mathrm{barrier}}$, hence $p_{\mathrm{fix}}$ agrees with Kimura’s formula for $s_{\mathrm{inv}}>s_{c}$.
The above argument also explains the behavior of mean absorption/extinction time. It grows and decays exponentially in the region $\sqrt{4\eta/N_{\mathrm{eff}}}<s_{\mathrm{inv}}\lesssim\eta$ due to the required time to escape the coexistence region. More precisely, the time grows exponentially as $s_{\mathrm{inv}}$ increases and the coexistence region becomes farther from $f=0$. The growth continues till $s_{\mathrm{inv}}=s_{c}$. For $s_{\mathrm{inv}}>s_{c}$, the coexistence region becomes closer to fixation, and the mean absorption time decays exponentially as fixation becomes more likely. For the mean extinction time, although the time for escaping from the coexistence region to $f=0$ is even longer, it also becomes exponentially unlikely that the mutant becomes extinct after reaching the coexistence region. Namely, a mutant becomes extinct mainly due to the first barrier at $f=f_{\mathrm{barrier}}$. As a result, the mean extinction time reaches maximum at $s_{\mathrm{inv}}=s_{c}$ and decays exponentially for $s_{\mathrm{inv}}>s_{c}$.
Appendix H Mean extinction time
Using the above qualitative picture of coexistence region, we can derive analytical estimates for the mean extinction time of the mutant. It is also similar to the establishment time in population genetics, which is the time when selection dominates over drift and drives a mutant with positive $s$ to fixation. As explained in the main text, the extinction or establishment time is less sensitive to model details and is easier for analytical estimation.
For the regimes without the coexistence region, i.e. $s_{\mathrm{inv}}<\sqrt{4\eta/N_{\mathrm{eff}}}$ or $s_{\mathrm{inv}}>\eta$, the extinction time can be approximated by the time required to reach the barrier at $f=f_{R=1}$; see Eqs. (141) and (142). For $s_{\mathrm{inv}}<\sqrt{4\eta/N_{\mathrm{eff}}}$, the barrier is reflecting and the mutant becomes extinct in the same order of magnitude of time after reaching the barrier. For $s_{\mathrm{inv}}>\eta$, the barrier is absorbing and the mutant can no longer become extinct after passing through the barrier. Now, the time for reaching the barrier can be estimated using Eq. (134); since $\Delta t$ increases with $f$, the required time to reach $f=f_{R=1}$ is dominated by $\Delta t(f_{R=1})$. Therefore, for $s_{\mathrm{inv}}<\sqrt{4\eta/N_{\mathrm{eff}}}$, the mean extinction time $T$ is
$T\sim\frac{2}{|s_{\mathrm{inv}}|}\left(\sqrt{1+\frac{4\eta}{N_{\mathrm{eff}}s_{\mathrm{inv}}^{2}}}-\mathrm{sgn}(s_{\mathrm{inv}})\right)^{-1}\,$, (146)
and for $s_{\mathrm{inv}}>\eta$,
$T\sim\frac{2}{s_{\mathrm{inv}}}\left(\sqrt{1-\frac{4\eta}{N_{\mathrm{eff}}s_{\mathrm{inv}}^{2}}}+1\right)^{-1}\,.$ (147)
Both cases match the expectation from population genetics that $T\sim 1/|s_{\mathrm{inv}}|$ asymptotically.
On the other hand, the mean extinction time deviates from the expectation from population genetics when the coexistence region is present, i.e. when $\sqrt{4\eta/N_{\mathrm{eff}}}<s_{\mathrm{inv}}<\eta$. In this regime, the mean extinction time is dominated by the exponentially long time to escape the coexistence region driven by the stochastic dynamics. The scale of the escape time can be estimated by translating the stochastic dynamics into 1D diffusion in the presence of a potential $V(f)$. Since $f$ relaxes to near equilibrium around the deterministic fixed point $f^{*}\simeq\tilde{s}_{\mathrm{inv}}$ in the long term (but before being absorbed at $f=0,1$), the probability distribution $p(f)$ can be obtained from the Fokker-Planck equation:
$\frac{\partial p}{\partial t}$ $=-\frac{\partial}{\partial f}\left(f(1-f)(s_{\mathrm{inv}}-\eta f)p\right)+\frac{\partial^{2}}{\partial^{2}f}\left(\frac{f(1-f)}{2N_{\mathrm{eff}}}p\right)$ (148)
$=\frac{\partial}{\partial f}\left(\frac{dV}{df}f(1-f)p\right)+\frac{\partial^{2}}{\partial^{2}f}\left(k_{B}T_{\mathrm{eff}}f(1-f)p\right)=0\,$, (149)
where we have used the language of statistical physics and defined a potential for $f$ (with minimum at $f=f^{*}$)
$V(f)=\frac{1}{2}\eta f^{2}-s_{\mathrm{inv}}f\,$, (150)
and an effective temperature
$k_{B}T_{\mathrm{eff}}=\frac{1}{2N_{\mathrm{eff}}}\propto D\,.$ (151)
As long as $f$ is not absorbed at 0 or $1$, the solution involves a Boltzmann distribution
$p(f)\propto\frac{e^{-V(f)/k_{B}T_{\mathrm{eff}}}}{f(1-f)}\,$, (152)
For large $\eta$ with $\alpha\gg 1$, we can focus on the exponential factor and approximate the escape time as $e^{\Delta V/k_{B}T_{\mathrm{eff}}}$ according to Kramer’s formula, where $\Delta V$ is the barrier height in the potential. In particular, the escape time to extinction (conditional on near equilibrium around $f^{*}$ and not escaping to fixation) is
$T_{0}\sim\exp\left(\frac{V(0)-V(\tilde{s}_{\mathrm{inv}})}{k_{B}T_{\mathrm{eff}}}\right)=\exp\left(\alpha^{2}\tilde{s}_{\mathrm{inv}}^{2}\right)\,.$ (153)
To further link $T_{0}$ to the mean extinction time $T$, we also need the probability that $f$ escapes to extinction instead of fixation, which is given by
$p_{0}\sim\frac{p(0)}{p(0)+p(1)}\sim\begin{cases}1&s<s_{c}\\ \frac{p(0)}{p(1)}=\exp\left(\alpha^{2}(1-\tilde{s}_{\mathrm{inv}})^{2}-\alpha^{2}\tilde{s}_{\mathrm{inv}}^{2}\right)&s>s_{c}\,.\end{cases}$ (154)
Therefore, we arrive at
$T\sim p_{0}T_{0}\sim\begin{cases}\exp\left(\alpha^{2}\tilde{s}_{\mathrm{inv}}^{2}\right)&s_{\mathrm{inv}}<s_{c}\\ \exp\left(\alpha^{2}(1-\tilde{s}_{\mathrm{inv}})^{2}\right)&s_{\mathrm{inv}}>s_{c}\,.\end{cases}$ (155)
In particular, the mean extinction time reaches maximum when $s_{\mathrm{inv}}=s_{c}\simeq\eta/2$, i.e.
$T_{\mathrm{max}}\sim\exp(\alpha^{2}/4)\,$, (156)
which is determined by $\alpha$ only.
Figure S6: Correlation between fixation probabilities and mean extinction times. In log scales, there is negative correlation between the fixation probability for neutral mutations and the maximum mean extinction time. Error bars denote standard errors from multiple instances of demographic noise and mutants.
It is interesting to observe that there are two independent quantities that are related to $e^{\alpha^{2}}$, namely the fixation probability for neutral mutations and the maximum mean extinction time. In particular, we expect that the maximum time is exponentially longer as the fixation probability is exponentially suppressed, since both exponential behavior is caused by parent-mutant coexistence as explained in the main text. Fig. S6 shows that the expectation is indeed realized in simulations.
Appendix I Methods, including details of simulations
We use simulations to compute the fixation probabilities and other statistical quantities within the full ecological models. The code for the below simulation and the figures in this paper can be found on https://github.com/Emergent-Behaviors-in-Biology/Pop-Gen-Ecology.
First, we sample the ecological properties of the community as described in Sec. A and B. We then evolve the community towards its unique steady state without demographic noise by solving the full differential equations using the LSODA method (with absolute and relative tolerances $10^{-12}$). The convergence is reached when all the derivatives $dN_{i}/dt,dR_{\alpha}/dt$ are less than $10^{-10}$. We can then use these results to compute the predicted values of $\eta$.
After that, we randomly choose one of the surviving species with equal probability as the parent. Since we only focus on a single mutation, the distribution for which parent is chosen does not affect the analysis. Each time, we introduce a mutant associated with the chosen parent, sampled as described in Sec. A and B. We then solve the full differential equations with demographic noise using a generalized Euler method as explained below. Due to the multiplicative nature of the noise, the simulation can become numerically unstable if we sample the noise directly. Instead, we make use of the result in [77]. Within a time step $dt$, we can approximate that the total growth rates of the species remain unchanged, since the growth rates involve averages over all the species or resource abundances. In other words, the species dynamics within time $dt$ becomes
$\frac{dN_{i}}{dt}=N_{i}g_{i}+\sqrt{2DN_{i}}\xi_{i}(t)\,$, (157)
where $g_{i}$ is the momentarily constant growth rate depending on the current species or resource abundances. It was noticed that the above equation can be solved exactly, and the solution is given by
$N_{i}(t+dt)\sim\frac{D(e^{g_{i}dt}-1)}{g_{i}}\cdot\mathrm{Gamma}\left(\mathrm{Poisson}\left(\frac{N_{i}(t)g_{i}}{D(1-e^{-g_{i}dt})}\right)\right)\,$, (158)
where $\mathrm{Gamma}(\alpha)$ is the gamma distribution with shape parameter $\alpha$ and unit scale. We define $\mathrm{Gamma}(0)$ as the Dirac delta distribution at 0. For each time $t$, we sample $N_{i}(t+dt)$ according to the above distribution. If resources are involved in the model, their abundances are updated at the same time using the ordinary Euler method. After that, to ensure uninvadability, we impose a hard wall for the abundances at $\lambda=10^{-7}$, i.e.
$N_{i}(t+dt)\rightarrow\max(\lambda,N_{i}(t+dt))\,$, (159)
and similarly for resources if present. We iterate the above process with a time step $dt=0.1$ till the parent or mutant abundance reaches $\lambda$, which is treated as extinction.
The parent and mutant trajectories in classic population genetics are simulated in the same way except that $g_{i}$ are now truly constants and the fitness difference is $s=g_{m}-g_{p}$.
To obtain fixation probabilities, we sample $10$ mutants with the same chosen parent. For each mutant, we adjust the value of $r_{m}$ (for generalized Lotka-Volterra model) or $m_{m}$ (for consumer-resource models) so that the invasion fitness is at a given value of $s_{\mathrm{inv}}$. We then run the above simulation $10^{3}$ times with different instances of demographic noise and count the number of fixations. We take the average of the above results to find the fixation probability $p_{\mathrm{fix}}(s_{\mathrm{inv}})$.
To obtain absoprtion and extinction times, we sample only one mutant for the chosen parent, then run the simulation $10^{3}$ times and record the time to fixation or extinction for each run.
Appendix J Parameters used in the figures
Throughout all figures, the initial mutant frequency is $f_{0}=0.09$. We set $A_{pm}=A_{mp}=\rho$ for all simulations of generalized Lotka-Volterra models.
In Fig. 1(f) and (h), we simulate the dynamics as in classic population genetics with the parent growth rate $g_{p}=0.1$, the parent abundance $N_{p}=5.0$, the fitness difference $s=0.01$, and the diffusion coefficient $D=0.5$. In Fig. 1(g) and (i), we simulate generalized Lotka-Volterra model with parameters
$S=50,\mu=2.0,\sigma=0.4,\mu_{r}=6.0,\sigma_{r}=0.6,\gamma=1,\rho=0.99,s_{\mathrm{inv}}=0.2,N_{p}=5.0,D=0.2\,.$ (160)
In all the subfigures in Fig. 2, we simulate generalized Lotka-Volterra model with parameters
$S=50,\mu=2.0,\mu_{r}=6.0,\sigma_{r}=0.6,\gamma=1\,.$ (161)
In Fig. 2(a) and (b), we further use the parameters
$\sigma=0.4,\rho=0.9,N_{p}=2.6$ (162)
We change the control parameter $\alpha$ by using $D=1.0$ in Fig. 2(a) and $D=0.04$ in Fig. 2(b). In Fig. 2(e), we use
$\sigma=0.4,s_{\mathrm{inv}}=0,N_{p}=4.8\,.$ (163)
In Fig. 2(f), the simulation parameters for the case of low packing are the same as those for Fig. 2(b). To control the extent of packing, we keep all the parameters unchanged except using $\sigma=0.7$ for the case with high packing.
In Fig. 3(a) and (b), we simulate generalized Lotka-Volterra model with parameters
$S=50,\mu=2.0,\sigma=0.4,\mu_{r}=6.0,\sigma_{r}=0.6,\gamma=1,\rho=0.9,N_{p}=2.0\,$, (164)
and $D=0.020,0.024,0.036$ corresponding to the three values of $N_{\mathrm{eff}}$. We focus on $D=0.024$ in Fig. 3(c)-(h). The selection-drift ratios in Fig.3 are calculated using Eq. (140) with the predicted values of $\eta$.
In Fig. 4(a), we simulate the dynamics as in classic population genetics with parameters
$s=0.14,N_{p}=2.0,D=0.01\,.$ (165)
In Fig. 4(b) and (c), we simulate generalized Lotka-Volterra model with parameters
$S=50,\mu=2.0,\sigma=0.4,\mu_{r}=6.0,\sigma_{r}=0.6,\gamma=1,\rho=0.9,N_{p}=2.0,D=0.01\,$, (166)
and $s_{\mathrm{inv}}=0.12,0.23$ in Fig. 4(b) and (c) respectively.
In Fig. S1(a), we simulate MacArthur consumer-resource model with parameters
$S=50,M=50,\mu=2.0,\sigma=0.8,\mu_{K}=5.0,\sigma_{K}=0.5,$ (167)
$\mu_{m}=2.0,\sigma_{m}=0.2,\kappa=1,\rho=0.9,N_{p}=5.8,D=0.1\,.$ (168)
In Fig. S1(b), we simulate the consumer-resource model with externally supplied resources with parameters
$S=50,M=50,\mu=2.0,\sigma=0.4,\mu_{K}=10.0,\sigma_{K}=0,$ (169)
$\mu_{m}=2.0,\sigma_{m}=0.2,\rho=0.9,N_{p}=42.0,D=0.2\,.$ (170)
In Fig. S2, the simulation parameters for the case of high packing (or higher $S^{*}/M^{*}$) are the same as in Fig. S1. To control the extent of packing, we keep all the parameters unchanged except using $S=20$ for the case with low packing (or lower $S^{*}/M^{*}$).
In Figs. S3 and S4, we simulate generalized Lotka-Volterra model with parameters
$S=50,\mu=2.0,\sigma=0.4,\mu_{r}=6.0,\sigma_{r}=0.6,\gamma=1,D=0.2\,.$ (171)
Fig. S3 uses $s_{\mathrm{inv}}=0$ and Fig. S4 uses $s_{\mathrm{inv}}=0,0.4,0.8$. To obtain the exponentially suppressed fixation probabilities, we repeat the simulation for each mutant $10^{4}$ times to count the number of fixations.
In Fig. S5, the selection-drift ratios are identical to the ones in Fig. 3.
In Fig. S6, the simulation parameters are the same as in Fig. 3(a) and (b), except that we now use $D=0.02,0.024,0.03,0.036,0.05$. To obtain the exponentially suppressed fixation probabilities, we repeat the simulation for each mutant $10^{5}$ times to count the number of fixations.